### v1

# Permutation-Invariant Spectral Learning via Dyson Diffusion

Abstract
Diffusion models are central to generative modeling and have been adapted to graphs by diffusing adjacency matrix representations. The challenge of having up to $n!$ such representations for graphs with $n$ nodes is only partially mitigated by using permutation-equivariant learning architectures. Despite their computational efficiency, existing graph diffusion models struggle to distinguish certain graph families, unless graph data are augmented with ad hoc features. This shortcoming stems from enforcing the inductive bias within the learning architecture. In this work, we leverage random matrix theory to analytically extract the spectral properties of the diffusion process, allowing us to push the inductive bias from the architecture into the dynamics. Building on this, we introduce the Dyson Diffusion Model, which employs Dyson’s Brownian Motion to capture the spectral dynamics of an Ornstein–Uhlenbeck process on the adjacency matrix while retaining all non-spectral information. We demonstrate that the Dyson Diffusion Model learns graph spectra accurately and outperforms existing graph diffusion models.
1 Introduction
Diffusion models are a key class of generative models based on noising data with Stochastic Differential Equations (SDEs) and learning their time reversal (Sohl-Dickstein et al., 2015; Song et al., 2021; Ho et al., 2020). They provide state-of-the-art results in many domains such as audio (Zhang et al., 2023) and vision (Croitoru et al., 2023). Generalizing diffusion models from Euclidean space to graphs offers promising applications in numerous areas, such as biology (Watson et al., 2023) or combinatorial optimization (Sun & Yang, 2023). However, while diffusing adjacency matrix representations is straightforward and popular (Niu et al., 2020; Jo et al., 2022; Vignac et al., 2022), this approach faces a major obstacle: Each graph with $n$ vertices has up to $n!$ representations as adjacency matrices. Therefore, if one aims to use a diffusion model on the space of matrices, one must learn $n!$ representations per graph. This is not feasible. Previous works tackled this problem by shifting the inductive bias of permutation invariance to the learning algorithm: If the neural network was permutation equivariant, training on one of the (up to $n!$ many) matrix representations would suffice. For example, Niu et al. (2020) and Jo et al. (2022) used message-passing graph neural networks (GNNs) while ConGress (Vignac et al., 2022) applied graph transformers. However, these learning architectures have a “blind spot” detailed below.
Theoretical Limitations.  The “blind spot” arises from the limited ability of these models to solve Graph Isomorphism ($\mathrm{GI}$): determining if two graphs are structurally identical regardless of node labeling. While permutation equivariance ensures that the model produces consistent outputs for different representations of the same graph, it does not guarantee that the model can differentiate between two structurally different (non-isomorphic) graphs. The failure is a result of how these architectures aggregate information: permutation-symmetric operations – message passing in GNNs and self-attention in graph transformers – can collapse distinct graphs with similar neighborhoods to the same representation, treating them as identical. More formally, $\mathrm{GI}$ is a challenging problem in algorithm theory, and it remains unknown whether $\mathrm{GI}\in\mathsf{P}$ (Babai, 2016). Since a polynomial time (learning) algorithm perfectly distinguishing all graphs would prove $\mathrm{GI}\in\mathsf{P}$, contemporary (polynomial-time) graph learning algorithms must compromise on expressivity.
Extracting Permutation-Invariant Information from Graph Diffusion. When diffusing an entire adjacency matrix, state-of-the-art work pushes the entire inductive bias into the learning algorithm (Niu et al., 2020; Jo et al., 2022) with possible data augmentation (Huang et al., 2022; Vignac et al., 2022; Xu et al., 2024). However, as we show below, this is neither necessary nor desirable (see the previous discussion on blind spots). Using techniques from random matrix theory, we show that an OU diffusion on the graph can be dissected into diffusion of the (permutation invariant) spectrum and diffusion of the (permutation-dependent) eigenvectors. Our method therefore allows us to learn the spectrum while preserving all remaining information. Moreover, since the spectrum is inherently permutation-invariant, we can parameterize the score using a much broader range of learning architectures, expanding the scope to architectures able to distinguish between graphs that are equivalent in the Weisfeiler-Leman (WL) sense (Morris et al., 2019; Xu et al., 2018). [^1] We give a theoretical discussion in form of the WL-equivalence class in Section 2 and demonstrate the challenge of those architectures empirically in Figure 2.
Figure 1: Dyson Diffusion model and its application to graph spectra: A graph on $n$ vertices (a) has up to $n!$ representations as adjacency matrices (b). For an OU-driven diffusion on any adjacency matrix, the permutation-invariant spectrum (c) evolves according to the same SDE (Dyson-BM). An exemplary path of the $n$, non-intersecting, eigenvalues is shown. The marginals of the invariant density for the $\lambda_{k}$ are depicted on the far right. The DyDM diffusion model learns the score $s(\lambda,t)$ (highlighted in yellow) to generate spectra via the time-reversed SDE (5).
Information in the Spectrum. The spectrum of a graph encodes key structural features including connectivity, expansion, and subgraph patterns. Moreover, non-isomorphic WL-equivalent graphs typically have distinct spectra (Huang & Yau, 2024). An idea, therefore, is to augment the graph data based on spectral information (Vignac et al., 2022; Xu et al., 2024). Our work is based on an entirely different method, exploiting analytical insight from random matrix theory to dissect the spectral from the remaining information, allowing to push the inductive bias from the architecture to the dynamics.
Dyson’s Brownian Motion. For an OU process on the space of symmetric matrices, the eigenvalues follow a well-characterized stochastic differential equation (SDE), the so-called Dyson Brownian Motion (DBM), which is inherently permutation invariant. Thus – in contrast to Niu et al. (2020); Jo et al. (2022); Vignac et al. (2022) – the score of DBM can be parameterized with any (not necessarily permutation invariant) neural network. Moreover, contrary to Luo et al. (2024), the remaining information is not lost (see Theorem 3.2).
Contributions. The main contributions of this work are as follows.
  • We introduce the novel Dyson Diffusion Model (DyDM) in Section 3, which extracts the spectral dynamics from an OU-driven diffusion. DyDM allows to learn the spectra of graphs without the need of GNNs or graph transformers while preserving the remaining information of the graph and allowing us to compute an eigenvector SDE.
  • We demonstrate in Section 4 that DyDM is more effective than existing GNN-based and graph-transformer-based methods for learning graph spectra.
  • We illustrate the struggle of GNN-based graph diffusion models in Figure 2 and Section 4.1.
Beyond graphs, our framework applies generally to symmetric matrices where spectral information is often key. For instance, in statistics, it can encode the importance of principal components (James et al., 2023), while in dynamical systems, [^2] For instance, Markov jump dynamics in detailed balance systems where the generator is symmetric in the steady state basis, see Pavliotis (2014). it reflects the stability and timescales of linear operators.
Notation. We work on the set of symmetric real matrices $\mathrm{Sym}(\mathbb{R}^{n\times n}):=\{A\in\mathbb{R}^{n\times n}:A^{T}=A\}$. The positive integers up to $n$ are denoted by $n]:=\{1,\ldots n\}$. We consider undirected graphs $G=(V,E,w)$ where $V$ is a finite set with edges $E\subseteq\{S\subseteq V:1\leq|S|\leq 2\}$ allowing for self-loops and weights $w:E\to\mathbb{R}$. The family of such graphs of size $n$ is $\mathcal{G}^{n}:=\{G=(V,E,w):|V|=n\}$. For a graph $G\in\mathcal{G}^{n}$ with $\mathbb{R}^{V}=\{f:V\to\mathbb{R}\}$ being the space of functions from $V$ to $\mathbb{R}$, we let $A\equiv A(G):\mathbb{R}^{V}\to\mathbb{R}^{V}$ be the adjacency operator, defined for $f\in\mathbb{R}^{V}$ and $v\in V$ as the weighted sum of $f$ applied to the neighbors of $v$ as
$Af(v)\coloneqq\sum_{\{u,v\}\in E}w\left(\{u,v\}\right)f(u).$
As the graphs are undirected, the adjacency operator is self-adjoint with respect to the standard inner product on $\mathbb{R}^{V}$. The operator $A$ therefore has $n$ real eigenvalues $\lambda_{1}\geq\lambda_{2}\geq\ldots\geq\lambda_{n}$. We denote the ordered spectrum of the adjacency operator by
$\lambda(G)\coloneqq\lambda(A(G))\coloneqq\{(\lambda_{1},\lambda_{2},\ldots,\lambda_{n})\,:\,\lambda_{1}\geq\ldots\geq\lambda_{n}\}.$
Unless defined explicitly otherwise, we use $n\in\mathbb{N}$ for the (vertex) size of the graph, and $N\in\mathbb{N}$ for the number of samples. We denote by $s\sim\mathcal{N}(0,I_{d})$ that the $d$-dimensional random vector $s$ has multivariate normal distribution with 0 mean and unit covariance $I_{d}$.
2 Limitations of Existing Diffusion Models
We consider the following matrix-valued Ornstein–Uhlenbeck (OU) SDE starting from some data $M(0)\in\mathrm{Sym}(\mathbb{R}^{n\times n})$ given for $1\leq i\leq j\leq n$ by
$\mathrm{d}M_{ji}(t)=\mathrm{d}M_{ij}(t)=-\beta M_{ij}(t)\mathrm{d}t+D_{ij}\mathrm{d}B_{ij}(t),$ (1)
with diffusion coefficient $D_{ij}\coloneqq\sqrt{(1+\delta_{ij})\alpha}$ for any constants $\alpha,\beta\in\mathbb{R}^{+}$, where $B_{ij}(t)=B_{ji}(t)$ are independent Brownian motions and $\delta_{ij}=1$ if and only if $i=j$, and 0 otherwise. We consider eq. (1) on the space of symmetric matrices to represent undirected graphs $G\in\mathcal{G}^{n}$.
Equation 1 is an entry-wise OU process that preserves the symmetry of the matrix. In standard diffusion models we run eq. 1 until some time $T>0$ from data samples. To fix notation, denote by $p_{t}$ the distribution of $M(t)$ induced by (1) for $t\geq 0$.
The time-reversal of the diffusion (1) over the time interval $[0,T]$ is a diffusion initialized by sampling $M(T)\sim p_{T}$ and satisfying for $1\leq i\leq j\leq n$
$\mathrm{d}M_{ij}(t)=-\left\{\beta M_{ij}(t)+D_{ij}^{2}\left[s(M(t),t)\right]_{ij}\right\}dt+D_{ij}\mathrm{d}\bar{B}_{ij}(t),$ (2)
for independent Brownian motions $\bar{B}_{ij}(t)$ (Anderson, 1982; Song et al., 2021). In Equation 2, $s(M,t)$ represents the score matrix at time $t\in(0,T]$, i.e., $s(M,t)\coloneqq\gradient_{M}\log p_{t}(M)$. This score is intractable but, as the OU process has tractable Gaussian transition densities, we can obtain an estimate $s_{\theta}(M,t)$ of it by minimizing the denoising score matching loss (Song et al., 2021)
$L(\theta)=\operatorname{\mathbb{E}}_{t\sim\mathrm{Unif}[0,T],M(0)\sim p_{0},M(t)\sim p_{t\mid 0}(\cdot|M(0))}\left[\norm{s_{\theta}(M(t),t)-\gradient_{M(t)}\log p_{t}(M(t)|M(0))}^{2}_{2}\right].$ (3)
Approximate samples of $p_{0}$ can then be obtained by simulating an approximation of Equation 2 obtained by sampling $M(T)$ from the Gaussian invariant distribution of (1), that is, $M_{ij}^{\rm inv}\sim\mathcal{N}(0,\alpha(1+\delta_{ij})/2\beta)$, and using $s_{\theta}(M,t)$ in place of $s(M,t)$.
Challenges posed by graphs. When working with graphs, to obtain an adjacency matrix for a given weighted graph $G=(V,E,w)\in\mathcal{G}^{n}$, we need to fix an ordering $(v_{1},\ldots,v_{n})$ of all vertices. In fact, given the ordering $(v_{1},\ldots,v_{n})$, the associated matrix $M=(m_{ij})_{1\leq i,j\leq n}\in\mathrm{Sym}(\mathbb{R}^{n\times n})$ has entries $m_{ij}=w(\{v_{i},v_{j}\})$ if $\{v_{i},v_{j}\}\in E$ and 0 otherwise. The challenge is that the adjacency operator $A(G)$ for a given graph $G\in\mathcal{G}^{n}$ admits up to $n!$ distinct adjacency matrices. For instance, in Figure 1 we see four different matrix representations of the same graph.
Why should we enforce this inductive bias? One could think that the reason for using the inductive bias stems from wanting that graph generative models assign uniform probability to each of the (up to) $n!$ many representations. But this can be easily achieved by applying a random independent permutation to the output of the generative model.
Instead, the challenge stems from the following problem: In diffusion models, we learn a function (i.e. the score) on a set of graphs, say [^3] Importantly, $\Omega$ refers to the set of graphs and not the set of (permutation-sensitive) adjacency matrices. $\Omega$, rather than the distribution directly. Learning on the adjacency representations would correspond to learning on $\Omega\times S_{n}$. We demonstrate on a toy example in Corollary K.1 that not leveraging the inductive bias, i.e. learning on a space of size $\Omega\times S_{n}$, leads to an explosion of the average mean squared error: Learning a function on a fixed number of $k$ objects (say unweighted graphs on $n$ nodes) from $N$ samples leads to a mean squared error of order $\Theta(n!/N)$. In contrast, making use of the inductive bias leads to an average mean squared error of $\Theta(1/N)$. Noting that already on $n=10$ nodes, we have $n!>3\cdot 10^{6}$, we see the clear argument for using the inductive bias.
Figure 2: Struggle of GNN-based and graph-transformer-based models with two WL-equivalent graphs: Graphs A and Graphs B are WL-equivalent, but non-isomorphic. Also physically, they have very different properties, such as different temperature factors (c) and a different cut size. Upon training on a $80\%$ Graph A and $20\%$ Graph B dataset, state-of-the-art GNN-based (EDP-GNN,GDSS) and graph-transformer-based (ConGress) models learn the WL-equivalence class quickly but fail to generate the underlying distribution among the two graphs, with some even predominantly hallucinating WL-equivalent but non-isomorphic graphs (d).
Cost of pushing inductive bias entirely into architecture. One solution would be to impose the inductive bias in the learning architecture. This is what state-of-the-art graph diffusion models do (Niu et al., 2020; Jo et al., 2022; Vignac et al., 2022).
However, using these architectures comes at a cost. As argued in the introduction, since $\mathrm{GI}\in\mathsf{P}$ remains unknown, we expect some limitations. In the case of GNNs this compromise has been precisely characterized: GNNs cannot distinguish between the large families of so-called Weisfeiler-Leman (WL) equivalent graphs (Morris et al., 2019; Xu et al., 2018). For example, all $k$-regular graphs on $n$ vertices for any fixed $k\in\mathbb{N}$ are WL-equivalent (see Lemma 2.1 below and its proof in Appendix L) and therefore indistinguishable for these architectures, which may lead to hallucination.
Lemma 2.1 (WL-equivalence of $k$-regular graphs).
For every fixed $n,k\in\mathbb{N}$, all $k$-regular graphs $G\in\mathcal{G}^{n}$ are WL equivalent. Moreover, every graph $G\in\mathcal{G}^{n}$ that is WL equivalent to a $k$-regular graph is $k$-regular.
This is a vast class, since e.g. on $n=20$ vertices, there are $510^{\prime}489$ many non-isomorphic, connected 3-regular and thereby WL-equivalent graphs (Meringer, 1999). In particular, we demonstrate on a simple example in Figure 2 that both GNN- and graph-transformer-based methods fail to learn a distribution on two particular WL-equivalent graphs: during sampling, they either fail to learn the distribution on both WL-equivalent graphs or hallucinate other, WL-equivalent but non-isomorphic, graphs (Fig. 2d). This can also be seen during learning: The EDP-GNN model learns the WL-equivalence class quickly (after 500 epochs) but then keeps hallucinating non-isomorphic but WL-equivalent graphs, and in particular struggles to learn the right distribution between graphs $A$ and $B$ for the remaining 4500 epochs (see Appendix O for details). Importantly, those graphs are very different. For instance, if the graphs represented Gaussian Network Models for macromolecules (Tirion, 1996), physical observables such as the temperature factors in X-ray scattering (Haliloglu et al., 1997) would be clearly distinct, see Fig. 2c. Therefore, a diffusion model based on GNNs will suffer from this expressivity blind spot. More generally, any diffusion model relying on a graph-specific learning algorithm will have limited expressivity.
3 Dyson Diffusion Model
3.1 Dyson Brownian Motion
Dyson (1962) showed that the spectrum of eq. 1 behaves as $n$ positively charged particles in a one-dimensional Coulomb gas. These particles exhibit Brownian motion, but with a pairwise repulsion force proportional to their inverse distance so that their paths do not cross (see Fig. 1c). More precisely, Dyson proved that the spectrum of the entry-wise Ornstein-Uhlenbeck process from eq. 1 follows the SDE given in Theorem 3.1. Without loss of generality, we restrict the domain to the Weyl Chamber $C_{n}:=\left\{\lambda\in\mathbb{R}^{n}:\lambda_{1}>\ldots>\lambda_{n}\right\}.$
Theorem 3.1 (Eigenvalue SDE, Dyson (1962)).
Denote by $\lambda(t)=(\lambda_{1}(t),\ldots,\lambda_{n}(t))$ the ordered spectrum of the matrix-valued Ornstein-Uhlenbeck process $M(t)$ of eq. 1. Then assuming that the initial matrix $M(0)$ has simple spectrum, $\lambda(t)$ satisfies for all $1\leq k\leq n$ the stochastic differential equation
$\mathrm{d}\lambda_{k}(t)=\left(\alpha\sum_{\ell\neq k}\frac{1}{\lambda_{k}(t)-\lambda_{\ell}(t)}-\beta\lambda_{k}(t)\right)\mathrm{d}t+\sqrt{2\alpha}\mathrm{d}W_{k}(t),$ (Dyson-BM)
for $W_{1},\ldots,W_{n}$ independent standard Brownian motions. Moreover the unique stationary distribution of (Dyson-BM) has density
$p_{\mathrm{inv}}(\lambda)=\frac{1}{Z}\exp(-U(\lambda))\quad\quad\text{ for }\quad\quad U(\lambda)=\frac{\beta}{2\alpha}\sum_{k}\lambda_{k}^{2}-\sum_{k<\ell}\ln|\lambda_{k}-\lambda_{\ell}|,$ (4)
for $\lambda\in C_{n}$ and $Z$ a normalizing constant so that $p_{\mathrm{inv}}$ corresponds to a probability measure.
For completeness, a full proof of Theorem 3.1 is given in Appendix B, where we generalize a well-known proof to arbitrary coefficients $\alpha,\beta$. We note that the assumption of Theorem 3.1 that $M_{0}$ has simple spectrum is minor. Indeed, generic random graphs or matrices have simple spectra, and in the case of eigenvalues with higher multiplicity, we can perturb the spectrum to be simple.
From Dyson-BM we see that the eigenvalues perform a Brownian motion in a confining potential with a repulsion force: once a pair of eigenvalues $\lambda_{k},\lambda_{l}$ comes too close, they become repelled with a force $\alpha/(\lambda_{k}-\lambda_{l})$ inversely proportional to their separation. A remarkable property of Theorem 3.1 is that the evolution of the spectrum is decoupled from all other information about the matrix: the spectral SDE (Theorem 3.1) is independent of the eigenvectors. This is the key analytical insight that motivates our approach.
Furthermore, conditioned on the eigenvalues, the remaining information captured in form of the eigenvectors can be deduced as we show in Theorem 3.2. This generalizes a statement of Allez et al. (2014), and we give a proof in Appendix C.
Theorem 3.2 (Eigenvector SDE).
Denote by $(v_{1}(t),\ldots,v_{n}(t))$ the orthonormal eigenvectors associated to the eigenvalues of Theorem 3.1. Assuming that the initial matrix $M(0)$ has simple spectrum, $v_{k}(t)$ satisfies for $k\in[n]$ the stochastic differential equation
$\mathrm{d}v_{k}(t)=-\frac{\alpha}{2}\sum_{\ell\neq k}\frac{1}{(\lambda_{k}(t)-\lambda_{\ell}(t))^{2}}v_{k}(t)\mathrm{d}t+\sqrt{\alpha}\sum_{\ell\neq k}\frac{1}{\lambda_{k}(t)-\lambda_{\ell}(t)}v_{\ell}(t)\mathrm{d}w_{\ell k}(t)$ (Eigenvector-SDE)
for $\{w_{ij:i\neq j}\}$ standard Brownian motions independent of the eigenvalue trajectories, with $w_{ji}=w_{ij}$.
3.2 From the Dyson SDE to a Diffusion Model
Despite its advantages, constructing a diffusion model based on Dyson-BM poses several challenges, e.g. dealing with a singular drift, non-Gaussian conditional density, etc. (see Appendix J for details). As described below, with DyDM we overcome these obstacles and design an efficient diffusion model for the spectrum, which can distinguish between spectra of graphs that GNNs are blind to (Fig. 2) and which does not require ad hoc data augmentation.
The time-reversal of the Dyson-BM in the sense of Anderson (1982) reads
$\mathrm{d}\bar{\lambda}_{k}(t)=\left[\alpha\sum_{\ell\neq k}\frac{1}{\bar{\lambda}_{k}(t)-\bar{\lambda}_{\ell}(t)}-\beta\bar{\lambda}_{k}(t)-2\alpha[s(\bar{\lambda}(t),t)]_{k}\right]\mathrm{d}t+\sqrt{2\alpha}\mathrm{d}\overline{W}_{k}(t),$ (5)
where we aim to learn the score $s(\lambda,t)\coloneq\gradient_{\lambda}\log p_{t}(\lambda)$. Because the coefficients in Dyson-BM are non-Lipschitz, the applicability of Anderson (1982) is not immediate. Accordingly, in Appendix D we sketch a proof of existence and uniqueness of a strong solution and verify that Anderson’s time reversal applies.
Making the loss tractable. Learning the loss function $s(\lambda,t)$ as in eq. 3 is not feasible for the Dyson-BM, since a closed form of the conditional distribution $p_{t\mid 0}$ is not known in contrast to the OU process. To overcome this, we follow a derivation in the style of De Bortoli et al. (2022) to obtain the loss function – up to constants in $\theta$ – for any $h\in\mathbb{R}^{+}$
$\tilde{L}(\theta)=\operatorname{\mathbb{E}}_{t\sim\mathrm{Unif}[0,T],\lambda_{t}\sim p_{t},\lambda_{t+h}\sim p_{t+h\mid t}(\cdot|\lambda_{t})}\left[\norm{s_{\theta}(\lambda_{t+h},t+h)-\gradient_{\lambda_{t+h}}\log p_{t+h\mid t}(\lambda_{t+h}|\lambda_{t})}^{2}_{2}\right],$ (6)
where we will approximate the intractable $p_{t+h\mid t}$ with the Gaussian transition of the Euler-Maruyama approximation (see Appendix A) for details.
Handling singularities. Numerical solutions of Dyson-BM with a fixed step size are not practical, since Dyson-BM is singular at the boundary of the Weyl Chamber. A fixed step size leads to inaccuracies at the boundaries and may overshoot the singularity, leaving the Weyl Chamber. To overcome this, we implement an adaptive step-size algorithm which conditions on an event of probability $1$ (non-crossing) and hence does not change marginal densities. [^4] Note that the adaptive step-size is only used in the forward simulation, whereas the objective is evaluated always on a fixed grid to ensure correctness. The step-size controller is described in Algorithm 2 and in Section F.1 in detail.
Schedule. Dyson’s conjecture states that for $\alpha=\frac{1}{n}$ and $\beta=\frac{1}{2}$ Dyson-BM converges to the invariant distribution in time $\Theta(1)$, while the majority of eigenvalues mixes locally already in time $\Theta(1/n)$ (Yang, 2022). Hence, the fine-grained structure will be mixed in the time interval $\left(0,\Theta(1/n)\right)$. We show in Appendix E through time change that this conjecture can be applied to any choice of coefficients $\alpha,\beta$ as. It is thus sensible to choose an exponential schedule. We specify the particular choices in Section I.1.
Algorithm 1 DyDM training
1:Input: spectral samples $\lambda^{(1)},\ldots,\lambda^{(N)}\in\mathbb{R}^{n}$, schedule $\mathcal{T}=\{t_{j}\}$
2:for each sample $i\in[N]$ do
3:  $t\leftarrow 0$
4:  $\lambda(t)\leftarrow\lambda^{(i)}$
5:  while $t<T$ do $\triangleright$ Diffuse sample
6:   Let $u\sim\mathcal{N}(0,I_{n})$
7:   $\delta t\leftarrow\text{ForwardStepsizeController}(\lambda(t),u)$ $\triangleright$ Conditions on non-intersection, see Appendix F.1.
8:   if $\delta t<\delta t_{\text{min}}$ then
9:     Skip step $\triangleright$ Rare event: Skip tiny steps
10:   end if
11:   $\delta t\leftarrow\min\{\delta_{t},t_{\text{fix}}-t\}$ where $t_{\text{fix}}=\min\{t_{j}\in\mathcal{T}:t_{j}>t\}$
12:   $\lambda(t+\delta t)\leftarrow$ Euler-Maruyama step of Dyson-BM with step size $\delta t$ and noise $u$
13:   $t\leftarrow t+\delta t$
14:  end while
15:end for
16:Update $s_{\theta}$ using loss $\tilde{L}(\theta)$ along paths $\lambda^{(1)}(t),\ldots,\lambda^{(N)}(t)$ on schedule $\mathcal{T}$ using eq. 18
17:Output: $s_{\theta}$.
Figure 3: Dyson Diffusion Model (training): The Dyson-BM is evolved forward in time with an adaptive step size ensuring that the paths remain in the Weyl Chamber. The step-size controller conditions on the probability 11 event of non-crossing as detailed in Section F.1 and Algorithm 2.
Equilibrium shooting mechanism. During inference, we require access to the learned score. In the forward diffusion, solely numerical errors due to the time discretisation may lead to crossings of singularities. On the one hand, when going backwards in time, due to inconsistencies in the learned score, the repulsion in eq. 5 might be too weak and the sample path may leave the Weyl Chamber for any sensible step size. Since the score is not defined outside the Weyl Chamber, this is problematic. On the other hand, the step size obtained by conditioning on not leaving the Weyl Chamber in this ill-trained point would be so small that the numerical solver would get stuck. To overcome this, we incorporate a shooting mechanism: If the repulsion force is too weak to prevent crossing of the singularity, resulting in a microscopic step size upon conditioning to remain in the Weyl Chamber, we repel with the invariant-state drift, i.e., we replace the learned score with the score in the invariant state (see Appendix G). This mechanism ensures that we stay in the Weyl Chamber, while minimizing its impact on the distribution. With well-tuned parameters, the shooting gets rarely triggered (less than $0.5\%$ of steps) but is essential, as already a single event would cause getting stuck (upon conditioning) or leaving the Weyl Chamber.
Sampling from Invariant Distribution To sample from the invariant distribution of $\lambda(t)$ given by eq. 4 we exploit the connection between $\lambda(t)$ and $M(t)$: We first sample from the Gaussian invariant distribution of $M(t)$ and then perform an eigendecomposition.
Comparison to direct simulation. One may wonder why we do not sample from the matrix-valued OU process at any time $t$ directly, then perform an eigendecomposition to get $\lambda(t)$, determine $\lambda(t+dt)$, and learn the score network from the increment using eq. 6. This way, learning would be simulation-free and the Dyson SDE would only be needed for the derivation of the loss function as well as the backwards dynamics. The reason for not pursuing this is efficiency and precision. We found that the above way takes $150$ times longer for graphs of size $n=10$, since (accurate) eigendecomposition is computationally costly. Doing so at every step explodes the costs. Moreover, note that we train on the sample of increments, so that from a simulation until time $t$, we learn on the entire generated path. Through our efficient implementation of the SDE, the learning process has running time (up to smaller time steps performed by the adaptive step size controller) on the order of a simulation-free diffusion model.
4 Experiments
We empirically evaluate DyDM against several state-of-the-art graph diffusion models. First, we compare them on a simple bimodal distribution between two WL-equivalent graphs, illustrating the struggle of purely GNN-based and graph-transformer-based methods (Section 4.1). Next, we carry out a comparison on a standard graph benchmark datasets (Section 4.1) and demonstrate scalability on a larger dataset ($15^{\prime}000$ graphs).
Table 1: Statistical distances of DyDM compared to standard models: DyDM learns the spectrum better in both the $n$-dimensional mean and the marginal Wasserstein sense, without requiring ad hoc data augmentation. Results are rounded to two decimal places, and results ()(*) are equal until the fourth decimal place, and ()(**) until the third decimal place. Exact numbers are provided in Section N.6.
Dataset
WL-Bimodal
Community Small Brain
Distance $\mu$ $\mathcal{W}_{\mathrm{marg}}$ $\mu$ $\mathcal{W}_{\mathrm{marg}}$ $\mu$ $\mathcal{W}_{\mathrm{marg}}$
DyDM (ours) $\bm{0.02}$ $\bm{0.01^{*}}$ $\bm{0.07}$ $\bm{0.02}$ $\bm{0.05}$ $\bm{0.03^{**}}$
EDP-GNN 0.13 0.08 0.42 0.14 - -
GDSS 0.23 0.13 0.42 0.14 - -
ConGress 0.38 0.16 0.27 0.11 - -
DiGress (no trick) 1.06 0.29 2.51 0.45 - -
DiGress (trick) 0.03 $\bm{0.01^{*}}$ 0.09 0.03 0.12 $\bm{0.03^{**}}$
4.1 Methodology
We compare to the GNN-based models EDP-GNN (Niu et al., 2020) and GDSS (Jo et al., 2022), as well as graph-transformer-based ConGress and DiGress (Vignac et al., 2022), where among all the models only DiGress uses a data-augmentation trick: it adds certain graph features, including cycle counts and the first $6$ eigenvalues (for details, see Section 5), but this trick is supposedly inessential for building a good model (Vignac et al., 2022). The trick, however, improves the learning of certain features, but not necessarily other subgraph structures (Wang et al., 2025). We thus compare to both the Digress model “without the trick”, i.e., just a graph (Markov chain) diffusion model, and with the data-augmentation trick.
As we evaluated the full spectrum and the published experimental results only showed partial or no information about the spectrum, we need explicit access to the samples. Since not all models (Vignac et al., 2022; Niu et al., 2020) report snapshots, we had to retrain those also on the standard datasets. We report our code and all the samples of all models in Github  [^5] See https://github.com/schwarzTass/DyDM-code.
WL-Bimodal. Here we train the simple bimodal distribution in Fig. 2, consisting of $80\%$ graph A and $20\%$ graph B, which are WL equivalent. The dataset has $N=5^{\prime}000$ random permutations of A and B, and we follow the standard test/train split procedure (Jo et al., 2022; You et al., 2018; Niu et al., 2020) using $80\%$ of the data as train data and the remaining $20\%$ as test data. We performed hyperparameter tuning of the comparison models as described in Appendix N.
Community. Being a standard benchmark (Niu et al., 2020; Jo et al., 2022; You et al., 2018), we include it for comparison. However, due to heavy undersampling, we only test for memorization. From our perspective, memorization is the best that can be tested with said benchmark, and we elaborate on this in Section N.1
Brain. From the human connectome graph we drew $15^{\prime}000$ ego-graphs (i.e. the induced subgraph of neighborhoods) of size $5$ to $10$ vertices Amunts et al. (2013); Rossi & Ahmed (2015). We give details in Appendix M.
4.2 Results
In Table 1 we report mean distances and marginal Wasserstein distances of the spectra. Explicitly, if $\nu_{\mathrm{test}}$ is the distribution of the spectrum of the test dataset and $\nu_{\mathrm{samp}}$ the distribution of our samples, then the distance between the means in $\mathbb{R}^{n}$ is $\mu(\nu_{\mathrm{samp}},\nu_{\mathrm{test}})=\norm{\mathbb{E}_{\lambda\sim\nu_{\mathrm{samp}}}[\lambda]-\mathbb{E}_{\lambda\sim\nu_{\mathrm{test}}}[\lambda]}_{2}.$ For statistical feasibility, instead of calculating the full Wasserstein distance we use the averaged marginal Wasserstein distance given by $\mathcal{W}_{\mathrm{marg}}(\nu_{\mathrm{samp}},\nu_{\mathrm{test}})=\frac{1}{n}\sum_{k=1}^{n}\mathcal{W}((\nu_{\mathrm{samp}})_{k},(\nu_{\mathrm{test}})_{k}),$ for $(\nu_{\mathrm{samp}})_{k}$ the marginal distribution in dimension $k$ and $\mathcal{W}$ the Wasserstein distance between two one-dimensional distributions. Using these metrics we can evaluate both marginal and high-dimensional effects.
These metrics also reveal the limitations of GNN- and Graph-Transformer-based models on the simple WL-Bimodal dataset as described in Figure 2, we also see that this extends to the real-world benchmark dataset Community Small. DyDM, on the other hand, consistently overcomes these issues. Even if we compare to the model with ad hoc feature augmentation (Digress with “trick”, on which we elaborate in Section 5), DyDM – which does not employ feature augmentation – either improves on or is on par with the feature augmented DiGress. We demonstrate that this performance still holds when working on large datasets, such as the $15^{\prime}000$ ego-graphs from the Brain dataset.
5 Related Work
The spectrum carries key features of graphs, and as such has been considered before, however, only as either auxiliary features for data augmentation (Vignac et al., 2022), or in a way which makes any remaining $\Theta(n^{2})$ degrees of freedom inaccessible, as we outline below.
Data augmentation. Vignac et al. (2022) acknowledge the importance of spectra of graphs and add plenty of auxiliary features, among them the first $6$ eigenvalues of the Graph Laplacian as graph-level features and the first $2$ non-zero eigenvectors as vertex-level features. This engineering trick is indeed helpful, yet it is a trick; as we show in Table 1 these auxiliary features are necessary for DiGress to provide a good model. This is in stark contrast to DyDM, where we build on an analytical expression of the evolution of all eigenvalues during diffusion of a symmetric matrix.
Spectral OU method. The use of spectral information for a graph diffusion model has been explored in recent work by Luo et al. (2024), which, however, performs an OU diffusion solely on the spectrum as if the eigenvectors were not impacted by diffusion, hence making any recovery of remaining information in the form of the eigenvectors impossible. The authors argue that the spectrum contains much information, so that upon sampling the spectrum from the OU-based model, the eigenvectors are simply sampled by taking the eigenvectors of a training sample chosen uniformly at random. However, this strategy is very limited, in that it allows only for a diffusion in $n$ parameters, losing $\Theta(n^{2})$ degrees of freedom irrecoverably. Since we learn in DyDM the spectrum of an OU diffusion on the entire graph, the remaining information remains accessible, see Theorem 3.2. This allows learning the eigenvectors beyond sampling uniformly from training data (Luo et al., 2024).
6 Extensions
A key property of DyDM is that it learns the spectrum while retaining all remaining $\Theta(n^{2})$ degrees of freedom accessible through Eigenvector-SDE. Since Dyson-BM decouples from Eigenvector-SDE, future work could implement a model of Eigenvector-SDE and generate entire graphs. To parameterize Dyson-BM, a graph-transformer model such as Jo et al. (2022) could be used; since Eigenvector-SDE is conditioned on the eigenvalue path which we learn without WL-blindness, we assume that this biases paths sufficiently far from each other to mitigate any WL-problems in classical graph-transformer based methods. Moreover, instead of learning the spectrum of the adjacency matrix $\lambda(A(G))$, one could also learn the combinatorial graph Laplacian $\lambda(L(G))$. Finally, we considered graphs with real (or integer) valued weights, while Dyson-BM is well-defined in other algebras. Hence, generalizations to complex-weighted graphs (Tian & Lambiotte, 2024) are straightforward: Instead of working on $\mathrm{Sym}(\mathbb{R}^{n\times n})$, one works on Hermitian ensembles. Similarly, matrices over the algebra of real quaternions may be considered (Dyson, 1962).
Beyond graphs. Since the Dyson SDE is defined on the domain of $\mathrm{Sym}(\mathbb{R}^{n\times n})$ (and even Hermitian matrices), it could be applied beyond graphs to other data. For instance, if correlations between $n$ points is measured in form of covariance matrices, DyDM could learn the spectrum and thereby quantify how strong the data clusters into (low-dimensional) principal components (Chen et al., 2015; Estavoyer & François, 2022; Hess, 2000).
7 Conclusion
Leveraging the analytical insights offered by Dyson’s Brownian Motion, we introduced DyDM, a diffusion model for spectral learning. Using techniques from Random Matrix Theory, we derived an evolution equation for the corresponding eigenvectors, which renders the remaining information about the underlying matrix available. On the domain of graphs, we demonstrated the struggle of existing learning architectures, e.g., GNN- and graph-transformer-based models. Building on the analytical insights, we decompose the dynamics, such that the spectral part is not constrained by inductive bias, thereby expanding the scope of suitable learning architectures. This way, DyDM can learn the spectrum (even of challenging graph families) without constraining to permutation-equivariant networks. This eliminates the hallucination, so that DyDM learns the distributions without struggle and without any need for data augmentation, as demonstrated experimentally. We hope that this approach opens a new direction in enforcing the inductive bias beyond the learning architecture in graph diffusion models.
Acknowledgments
We thank Jakiw Pidstrigach and Marc Roth for helpful discussions. Financial support from the European Research Council (ERC) under the European Union’s Horizon Europe research and innovation program (Grant Agreement No. 101086182 to A. G.), the Rhodes Trust (to T. S.) and the Swiss National Science Foundation (to C.K. grant number 235409) are acknowledged. This work used the Scientific Compute Cluster at GWDG, the joint data center of Max Planck Society for the Advancement of Science (MPG) and University of Göttingen. In part funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 405797229.
Appendix
The Appendix is structured as follows. We first derive the tractable loss $\tilde{L}(\theta)$ in Appendix A. We then provide proofs of the Equation Dyson-BM in Appendix B as well as for the Eigenvector-SDE in Appendix C. We then explain the applicability of Anderson’s time reversal (Appendix D), and provide a time-rescaling of Dyson-BM in Appendix E. In Appendix F, we provide the adaptive step size controller, followed by an explanation of the shooting mechanism (Appendix G), sampling procedure (Appendix H), and engineering details (Appendix I). We explain the challenges of Dyson’s Brownian Motion which we overcame with DyDM in Appendix J. We present our theoretical argument for using inductive bias in Appendix K and prove the Lemma on WL equivalence of $k$-regular graphs in Appendix L. For the experiments, we first elaborate on the datasets (Appendix M) followed by an explanation of our extensive benchmarking in Appendix N, including a comment about undersampling in some benchmark datasets in Section N.1. We show the learning dynamics of $4$ different runs of EDP-GNN in Appendix O.
Appendix A Making the loss tractable
In this section, we deduce a general loss formula for an SDE on $\mathbb{R}^{d}$ that will be applied to (Dyson-BM). We consider the process $X=(X(t))_{t\geq 0}$ determined by the SDE
$\mathrm{d}X(t)=a(t,X(t))\mathrm{d}t+b(t,X(t))\mathrm{d}W_{t}$
with initial condition $X(0)$. We assume throughout this section that $X(t)$ is absolutely continuous for all $t\geq 0$ and denote by $p_{t}$ the density of $X(t)$. We furthermore assume that the joint densities of $X(t)$ and $X(s)$ also have density for all $t,s\geq 0$ that we write as $p_{t,s}$. Observe that by Bayes formula for $t\geq s\geq 0$ we have for $x,y\in\mathbb{R}^{d}$ that
$p_{t,s}(y,x)=p_{t|s}(y|x)p_{s}(x),$ (7)
where $p_{t|s}(y|x)$ is the conditional density of $y$ given $x$.
The canonical loss arising from $s(y,t)=\gradient_{M(t)}\log p_{t}(M)$ is
$L^{\prime}(\theta)=\operatorname{\mathbb{E}}_{t\sim\mathrm{Unif}[0,T],M(0)\sim p_{0},M(t)\sim p_{t\mid 0}(\cdot|M(0))}\left[\norm{s_{\theta}(M(t),t)-\gradient_{M(t)}\log p_{t}(M(t))}^{2}_{2}\right].$ (8)
It is a well-known fact that $L^{\prime}(\theta)=L(\theta)+\mathrm{const}(\theta)$. We will first explain how the loss $L^{\prime}(\theta)$ from (8) and $\tilde{L}(\theta)$ from (6) are the same up to a constant, that is $L^{\prime}(\theta)=\tilde{L}(\theta)+\mathrm{const}(\theta)$, which therefore results in the same gradient descent as with $L(\theta)$.
We generalize the loss from (8) to weighing the time by a function $\eta$. So we consider the following generalized loss
$L(\theta)\coloneqq$ $\frac{1}{T}\int_{0}^{T}\eta(t)\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\norm{s_{\theta}(y,t)-\gradient_{y}\log(p_{t}(y))}_{2}^{2}\;p_{t,0}(y,x)\,\mathrm{d}y\mathrm{d}x\mathrm{d}t$
$=$ $\frac{1}{T}\int_{0}^{T}\eta(t)\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\norm{s_{\theta}(y,t)-\gradient_{y}\log(p_{t}(y))}_{2}^{2}\;p_{t\mid 0}(y\mid x)\mathrm{d}yp_{0}(x)\mathrm{d}x\mathrm{d}t,$ (9)
where $\eta(t)$ is some weighting function with $\int_{0}^{T}\eta(t)=1$. We wrote it in the above form with $p_{t\mid 0}$ denoting the conditional density, since the three integrals can be replaced by $\operatorname{\mathbb{E}}_{t\sim U{[0,T]}}\left[\cdots\operatorname{\mathbb{E}}_{x\sim p_{0}}\left[\operatorname{\mathbb{E}}_{y\sim p_{t\mid 0}(\cdot\mid x)}\left[\norm{\cdots}_{2}^{2}\right]\right]\right]$, which we can sample from if we assume (1) sample access to $p_{0}$, (2) known density at any $t$ given a dirac-delta $p_{0}$, (3) the term in the norm is tractable. (1) is assumed by the problem definition, (2) is known for an Ornstein-Uhlenbeck forward SDE, (3) can be solved by realizing that there is an equivalent loss function $\hat{L}(\theta)=L(\theta)+\rm const(\theta)$ where $p_{t}$ inside the norm is converted to a $p_{t\mid 0}$, which is known (gaussian density) in an OU setting.
Here, steps (2) and (3) fail. Note that by the polarization identity
$\norm{s_{\theta}(y,t)-\gradient_{y}\log(p_{t}(y))}_{2}^{2}=\norm{s_{\theta}(y,t)}_{2}^{2}+\norm{\gradient_{y}\log(p_{t}(y))}_{2}^{2}-2s_{\theta}(y,t)^{T}\gradient_{y}\log(p_{t}(y))$
We will rewrite the mixed term $s_{\theta}(y,t)^{T}\gradient_{y}\log(p_{t}(y))$ of the $L^{2}$ norm in such a way that the first quadratic term $\norm{s_{\theta}(y,t)}_{2}^{2}$ remains unchanged and the second quadratic term $\norm{\gradient_{y}\log(p_{t}(y))}_{2}^{2}$ is constant in $\theta$. First, we rewrite
$\frac{1}{T}\int_{0}^{T}\eta(t)\int_{\mathbb{R}^{n}\times\mathbb{R}^{n}}s_{\theta}(y,t)^{T}\gradient_{y}\log(p_{t}(y))p_{t,0}(y,x)\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}t$ (10)
$=$ $\frac{1}{T}\int_{0}^{T}\eta(t)\int_{\mathbb{R}^{n}}s_{\theta}(y,t)^{T}\gradient_{y}\log(p_{t}(y))p_{t}(y)\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}t$
$=$ $\frac{1}{T}\int_{0}^{T}\eta(t)\int_{\mathbb{R}^{n}}s_{\theta}(y,t)^{T}\gradient_{y}\left[p_{t}(y)\right]\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}t.$ (11)
We next observe that for $0\leq s^{\prime}\leq s$ we have the following:
$\gradient_{y}\left[p_{s}(y)\right]$ $=\gradient_{y}\left[\int_{\mathbb{R}^{n}}p_{s|s^{\prime}}(y|z)p_{s^{\prime}}(z)\mathop{}\!\mathrm{d}z\right]$
$=\int_{\mathbb{R}^{n}}\gradient_{y}p_{s|s^{\prime}}(y|z)p_{s^{\prime}}(z)\mathop{}\!\mathrm{d}z$
$=\int_{\mathbb{R}^{n}}\gradient_{y}p_{s|s^{\prime}}(y|z)\frac{p_{s,s^{\prime}}(y,z)}{p_{s|s^{\prime}}(y|z)}\mathop{}\!\mathrm{d}z$
$=\int_{\mathbb{R}^{n}}\gradient_{y}\log p_{s|s^{\prime}}(y|z)p_{s,s^{\prime}}(y,z)\mathop{}\!\mathrm{d}z.$
So we now perform in (11) for a small $h>0$ a change of variables to $t+h$ and apply the latter equality with $s=t+h$ and $s^{\prime}=t$. Then up to ignoring the boundary at 0 and $T$, and using by (7) that $p_{t+h,t}(y,z)=p_{t+h|t}(y|z)p_{t}(z)$ it follows that the loss from (6)
[…]
So it follows that
$\tilde{L}(\theta)=\frac{1}{T}\int_{0}^{T}\eta(t+h)\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\norm{s_{\theta}(y,t+h)-\gradient_{y}\log(p_{t+h|t}(y|z))}_{2}^{2}p_{t+h|t}(y|z)\mathop{}\!\mathrm{d}y\,p_{t}(z)\mathop{}\!\mathrm{d}z\mathop{}\!\mathrm{d}t$ (13)
is equal to $L(\theta)$ up to a constant term in $\theta$ (ignoring the boundary terms at 0 and $T$).
We will now make a series of approximations to calculate the loss $\tilde{L}(\theta)$. The first one is to approximate the integral $\int_{0}^{T}$ over $t$ by a sum $\sum_{i=1}^{k}$ over the time points $t_{0}<t_{1}<\ldots<t_{k}$ with $t_{0}=0$ and $t_{k}=T$. The second approximation we make is that we approximate the latter integral $\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}$ by sampling a path from $p_{t}$ at the time steps $t_{i}$. Indeed, we denote for each integer $1\leq r\leq N$ by $x^{(r)}_{t_{i}}$ the sample path of $p_{t}$. Moreover, we actually make the time grid also dependent on our sample path. So for each $1\leq r\leq N$, let $t_{0}^{(r)}<t_{1}^{(r)}<\ldots<t_{k^{(r)}}^{(r)}$ with $t_{0}^{(r)}=0$ and $t_{k^{(r)}}^{(r)}=T$ be the discretization of $[0,T]$. Thus, the overall loss can be approximated as
$\tilde{L}(\theta)\approx$ $\frac{1}{N}\sum_{r=1}^{N}\sum_{i=1}^{k^{(r)}}\frac{t_{i}^{(r)}-t_{i-1}^{(r)}}{T}\eta(t_{i}^{(r)})\norm{s_{\theta}(x_{i}^{(r)},t_{i}^{(r)})-\gradient_{y}\left[\log(p_{t_{i}^{(r)}\mid t_{i-1}^{(r)}}(x_{i}^{(r)}\mid x_{i-1}^{(r)}))\right]}_{2}^{2}$
$+\rm const(\theta).$ (14)
We finally approximate the incremental score function $\gradient_{x_{i}^{(r)}}\left[\log(p_{t_{i}^{(r)}\mid t_{i-1}^{(r)}}(x_{i}^{(r)}\mid x_{i-1}^{(r)}))\right]$ as follows. If $h>0$ is a small time step, we can approximate the conditional random variable $X_{t+h}\mid\mathcal{F}_{t}$ with $x:=X_{t}$ by
$X_{t+h}\mid\mathcal{F}_{t}$ $\approx x+a(t,x)h+b(t,x)N(0,h)$ (where $N(0,h)$ is a centered Gaussian RV with variance $h$)
$\sim\mathcal{\operatorname{\mathcal{N}}\left(,\,\right)}{x+a(t,x)h}{b^{2}(t,x)h},$ (15)
so that we have for the density
$p_{t+h|t}(y|x)$ $\approx\frac{1}{\sqrt{2\pi b^{2}(t,x)h}}\exp(-\frac{(y-x-a(t,x)h)^{2}}{2b^{2}(t,x)h}),$ (16)
which means for the scores
$\gradient_{y}\log\left(p_{t+h|t}(y|x)\right)$ $\approx-\frac{y-x-a(t,x)h}{b^{2}(t,x)h}.$ (17)
Thus, combining (A) and (17), the loss $\tilde{L}(\theta)$ can be approximated by the following as we use in our model:
$\frac{1}{N}\sum_{r=1}^{N}\sum_{i=1}^{k^{(r)}}\frac{t_{i}^{(r)}-t_{i-1}^{(r)}}{T}\eta(t_{i}^{(r)})\norm{s_{\theta}(x_{i}^{(r)},t_{i}^{(r)})-\frac{a(t_{i-1}^{(r)},x_{i-1}^{(r)})\cdot(t_{i}^{(r)}-t_{i-1}^{(r)})-(x_{i}^{(r)}-x_{i-1}^{(r)})}{b^{2}(t_{i-1}^{(r)},x_{i-1}^{(r)})\cdot(t_{i}^{(r)}-t_{i-1}^{(r)})}}_{2}^{2}$ (18)
Appendix B Spectral Dyson SDE
Theorem 3.1 (restated).
Denote by $\lambda(t)=(\lambda_{1}(t),\ldots,\lambda_{n}(t))$ the ordered spectrum of the matrix-valued Ornstein-Uhlenbeck process $M(t)$ of eq. 1. Then assuming that the initial matrix $M(0)$ has simple spectrum, $\lambda(t)$ satisfies for all $1\leq k\leq n$ the stochastic differential equation
$\mathrm{d}\lambda_{k}(t)=\left(\alpha\sum_{\ell\neq k}\frac{1}{\lambda_{k}(t)-\lambda_{\ell}(t)}-\beta\lambda_{k}(t)\right)\mathrm{d}t+\sqrt{2\alpha}\mathrm{d}W_{k}(t),$ (Dyson-BM)
for $W_{1},\ldots,W_{n}$ independent standard Brownian motions. Moreover the unique stationary distribution of (Dyson-BM) has density
$p_{\mathrm{inv}}(\lambda)=\frac{1}{Z}\exp(-U(\lambda))\quad\quad\text{ for }\quad\quad U(\lambda)=\frac{\beta}{2\alpha}\sum_{k}\lambda_{k}^{2}-\sum_{k<\ell}\ln|\lambda_{k}-\lambda_{\ell}|,$
for $\lambda\in C_{n}$ and $Z$ a normalizing constant so that $p_{\mathrm{inv}}$ corresponds to a probability measure.
Proof.
We prove the theorem in two parts.
Dyson SDE.
We first prove Dyson-BM. This part of the proof is based on Keating (2023) but generalizes it to arbitrary coefficients. For mathematical details on the $\alpha=\frac{1}{n}$, $\beta=0$ case, see also Anderson et al. (2009). Suppose $M(t)$ satisfies the SDE eq. 1. That is, with $M(0)\in\mathrm{Sym}(\mathbb{R}^{n\times n})$ we have
$\mathrm{d}M_{ij}(t)=-\beta M_{ij}\mathrm{d}t+D_{ij}\mathrm{d}B_{ij}$
with $D_{ij}\coloneqq\sqrt{(1+\delta_{ij})\alpha}$ for any constants $\alpha,\beta\in\mathbb{R}^{+}$ with $B_{ij}(t)=B_{ji}(t)\;\forall t>0$. Let $\lambda_{1}\geq\ldots\geq\lambda_{n}$ be the eigenvalues. We choose as $v_{1},\ldots,v_{n}$ an orthonormal basis of eigenvectors ($Mv_{k}=\lambda_{k}v_{k}$), which exists by the spectral theorem for symmetric matrices.
Due to symmetry, we may constrain the set of indices $(i,j)$ to $\mathcal{I}\coloneqq\{(i,j)\mid 1\leq i\leq j\leq n\}$. For any $k\in[n]$, the eigenvalue $\lambda_{k}$ can thus be seen as a function of the set of Itô processes $\{M_{\eta}\mid\eta\in\mathcal{I}\}$. Hence, we have by Itô’s Lemma,
$\mathrm{d}\lambda_{k}$ $=\underbrace{\frac{\partial\lambda_{k}}{\partial t}dt}_{\equiv 0}+\sum_{\eta\in\mathcal{I}}\frac{\partial\lambda_{k}}{\partial M_{\eta}}\mathrm{d}M_{\eta}+\frac{1}{2}\sum_{\eta,\xi\in\mathcal{I}}\frac{(\partial\lambda_{k})^{2}}{\partial M_{\eta}\partial M_{\xi}}\mathrm{d}M_{\eta}\mathrm{d}M_{\xi},$ (19)
where the first part is 0 since $\lambda_{k}(t)$ is only a function of the $M_{\eta}(t)$, not of time. By eq. 1, we get
$=\sum_{\eta\in\mathcal{I}}\left(-\beta M_{\eta}\frac{\partial\lambda_{k}}{\partial M_{\eta}}+\frac{1}{2}D_{\eta}^{2}\frac{\partial^{2}\lambda_{k}}{(\partial M_{\eta})^{2}}\right)\mathrm{d}t+D_{\eta}\frac{\partial\lambda_{k}}{\partial M_{\eta}}\mathrm{d}B_{\eta}.$ (20)
It remains to calculate the partial derivatives. In what follows, we successively apply properties of the spectrum in order to obtain equations for the partial derivatives. We have
$Mv_{k}$ $=\lambda_{k}v_{k}$ (21)
taking the partial derivative with respect to $M_{ij}$ for $(i,j)\in\mathcal{I}$ on both sides and applying the product rule yields
$\frac{\partial M}{\partial M_{ij}}v_{k}+M\frac{\partial v_{k}}{\partial M_{ij}}$ $=\frac{\partial\lambda_{k}}{\partial M_{ij}}v_{k}+\lambda_{k}\frac{\partial v_{k}}{\partial M_{ij}}.$ (22)
Further, by orthogonality of the eigenvectors we have for any $k,l\in[n]$
$v_{k}^{T}v_{l}$ $=\delta_{kl}$ (23)
$\frac{\partial v_{k}^{T}}{\partial M_{ij}}v_{l}+v_{k}^{T}\frac{\partial v_{l}}{\partial M_{ij}}$ $=0.$ (24)
which, taking the transpose, implies for $l=k$
$\frac{\partial v_{k}^{T}}{\partial M_{ij}}v_{k}$ $=v_{k}^{T}\frac{\partial v_{k}}{\partial M_{ij}}=0.$ (25)
Multiplying eq. 22 by $v_{l}^{T}$ from the left yields
$v_{l}^{T}\frac{\partial M}{\partial M_{ij}}v_{k}+\underbrace{v_{l}^{T}M\frac{\partial v_{k}}{\partial M_{ij}}}_{(I)}$ $=\underbrace{\frac{\partial\lambda_{k}}{\partial M_{ij}}v_{l}^{T}v_{k}}_{(II)}+\underbrace{\lambda_{k}v_{l}^{T}\frac{\partial v_{k}}{\partial M_{ij}}}_{(III)}$ (26)
where the terms simplify as follows.
$(I)$ $=(1-\delta_{kl})\lambda_{l}v_{l}^{T}\frac{\partial v_{k}}{\partial M_{ij}}$ (by eq. 25 )
$(II)$ $=\delta_{lk}\frac{\partial\lambda_{k}}{\partial M_{ij}}$ (27)
$(III)$ $=(1-\delta_{kl})\lambda_{k}v_{l}^{T}\frac{\partial v_{k}}{\partial M_{ij}},$ (by eq. 25 )
resulting in the following equation for $l=k$
$v_{k}^{T}\frac{\partial M}{\partial M_{ij}}v_{k}$ $=\frac{\partial\lambda_{k}}{\partial M_{ij}}$ (28)
and for $l\neq k$:
$v_{l}^{T}\frac{\partial M}{\partial M_{ij}}v_{k}+\lambda_{l}v_{l}^{T}\frac{\partial v_{k}}{\partial M_{ij}}$ $=\lambda_{k}v_{l}^{T}\frac{\partial v_{k}}{\partial M_{ij}}$ (29)
$v_{l}^{T}\frac{\partial M}{\partial M_{ij}}v_{k}$ $=\left(\lambda_{k}-\lambda_{l}\right)v_{l}^{T}\frac{\partial v_{k}}{\partial M_{ij}}.$ (30)
The matrix derivative is, trivially,
$\left(\frac{\partial M}{\partial M_{ij}}\right)_{kl}=\begin{cases}1&\text{for }(k,l)=(i,j)\text{ or }(l,k)=(i,j)\\ 0&\text{else}\end{cases},$ (31)
so that we can simplify eq. 28 to
$\frac{\partial\lambda_{k}}{\partial M_{ij}}$ $=(v_{k})_{i}(v_{k})_{j}(2-\delta_{ij})$ (32)
and eq. 30 to
$\left(\lambda_{k}-\lambda_{l}\right)v_{l}^{T}\frac{\partial v_{k}}{\partial M_{ij}}$ $=(v_{k})_{i}(v_{l})_{j}+(1-\delta_{ij})(v_{k})_{j}(v_{l})_{i}$ (33)
$=(v_{l})_{i}(v_{k})_{j}+(1-\delta_{ij})(v_{l})_{j}(v_{k})_{i}.$ (34)
Using that the $v_{k}$ are an ONB of eigenvectors, we can conclude with eq. 32 that the first summand of eq. 20 is
$\sum_{\eta\in\mathcal{I}}-\beta M_{\eta}\frac{\partial\lambda_{k}}{\partial M_{\eta}}$ $=-\beta\lambda_{k}.$ (35)
Since $\{v_{1},\ldots,v_{n}\}$ is an orthonormal basis, we may project on the basis vectors as follows
$\frac{\partial v_{k}}{\partial M_{ij}}$ $=\sum_{l\in[n]}v_{l}^{T}\frac{\partial v_{k}}{\partial M_{ij}}v_{l}$ (projection into ONB basis)
applying eq. 25 gives
$=\sum_{l\in[n]\setminus\{k\}}v_{l}^{T}\frac{\partial v_{k}}{\partial M_{ij}}v_{l}$ (36)
which yields with eq. 33
$=\sum_{l\in[n]\setminus\{k\}}\frac{1}{\lambda_{k}-\lambda_{l}}\left((v_{k})_{i}(v_{l})_{j}+(1-\delta_{ij})(v_{k})_{j}(v_{l})_{i}\right)v_{l}.$ (37)
To obtain the second order partial derivative of $\lambda_{k}$, we observe
$\frac{\partial^{2}\lambda_{k}}{\partial M_{ij}\partial M_{ij}}=$ $\frac{\partial}{\partial M_{ij}}\left((v_{k})_{i}(v_{k})_{j}(2-\delta_{ij})\right)$ (by eq. 32)
$=$ $(2-\delta_{ij})\left(\frac{\partial(v_{k})_{i}}{\partial M_{ij}}(v_{k})_{j}+(v_{k})_{i}\frac{\partial(v_{k})_{j}}{\partial M_{ij}}\right)$
$=$ $(2-\delta_{ij})\sum_{l\in[n]\setminus\{k\}}\frac{1}{\lambda_{k}-\lambda_{l}}\bigg((v_{l})_{i}(v_{k})_{j}(v_{l})_{i}(v_{k})_{j}+(v_{l})_{j}(v_{k})_{i}(1-\delta_{ij})(v_{l})_{i}(v_{k})_{j}$
$+(v_{l})_{i}(v_{k})_{j}(v_{l})_{j}(v_{k})_{i}+(v_{l})_{j}(v_{k})_{i}(1-\delta_{ij})(v_{l})_{j}(v_{k})_{i}\bigg)$ (by eq. 37)
$=$ $(2-\delta_{ij})\sum_{l\in[n]\setminus\{k\}}\frac{1}{\lambda_{k}-\lambda_{l}}\bigg((v_{l})_{i}^{2}(v_{k})_{j}^{2}+(v_{l})_{j}^{2}(v_{k})_{i}^{2}(1-\delta_{ij})$
$+(v_{l})_{i}(v_{l})_{j}(v_{k})_{i}(v_{k})_{j}(2-\delta_{ij})\bigg).$ (38)
These second order partial derivative are summed over $\mathcal{I}$ in eq. 20, so that combining eq. 38 with the definition of $D_{ij}$ yields
$\sum_{(i,j)\in\mathcal{I}}D_{ij}^{2}\frac{\partial^{2}\lambda_{k}}{(\partial M_{ij})^{2}}$ $=\sum_{j=1}^{n}\sum_{i=1}^{j}\alpha(1+\delta_{ij})(2-\delta_{ij})\sum_{l\in[n]\setminus\{k\}}\frac{1}{\lambda_{k}-\lambda_{l}}\bigg((v_{l})_{i}^{2}(v_{k})_{j}^{2}$
$+(v_{l})_{j}^{2}(v_{k})_{i}^{2}(1-\delta_{ij})+(v_{l})_{i}(v_{l})_{j}(v_{k})_{i}(v_{k})_{j}(2-\delta_{ij})\bigg)$ (39)
we reorder and note that since the summand is symmetric in $i,j$, we may change the range of summation and absorb the coefficient $2-\delta_{ij}$,
$=\alpha\sum_{l\in[n]\setminus\{k\}}\frac{1}{\lambda_{k}-\lambda_{l}}\sum_{j=1}^{n}\sum_{i=1}^{n}(1+\delta_{ij})\bigg((v_{l})_{i}^{2}(v_{k})_{j}^{2}$
$+(v_{l})_{j}^{2}(v_{k})_{i}^{2}(1-\delta_{ij})+(v_{l})_{i}(v_{l})_{j}(v_{k})_{i}(v_{k})_{j}(2-\delta_{ij})\bigg)$ (40)
where we realize that upon accounting for all $\delta_{ij}$, the first two summands are simply $\norm{v_{l}}_{2}^{2}\norm{v_{k}}_{2}^{2}$, while the third summand may we written an inner product
$=\alpha\sum_{l\in[n]\setminus\{k\}}\frac{1}{\lambda_{k}-\lambda_{l}}\left(2\norm{v_{l}}_{2}^{2}\norm{v_{k}}_{2}^{2}+2(\underbrace{v_{l}^{T}v_{k}}_{0})^{2}\right)$ (41)
since we chose an ONB, this simplifies to
$=2\alpha\sum_{l\in[n]\setminus\{k\}}\frac{1}{\lambda_{k}-\lambda_{l}}.$ (42)
It remains to determine an explicit expression of the Brownian motion in eq. 20. We have
$\sum_{\eta\in\mathcal{I}}D_{\eta}\frac{\partial\lambda_{k}}{\partial M_{\eta}}\mathrm{d}B_{\eta}$ $=\sum_{(i,j)\in\mathcal{I}}\sqrt{(1+\delta_{ij})\alpha}(2-\delta_{ij})(v_{k})_{i}(v_{k})_{j}\mathrm{d}B_{ij}$
using that $B_{ij}(t)=B_{ji}(t)$, we obtain
$=\sqrt{2\alpha}\sum_{i=1}^{n}\sum_{j=1}^{n}\sqrt{\frac{1+\delta_{ij}}{2}}(v_{k})_{i}(v_{k})_{j}\mathrm{d}B_{ij}$
we may now define $\mathrm{d}\tilde{B}_{k}$ as follows
$=\sqrt{2\alpha}\mathrm{d}\tilde{B}_{k}.$ (43)
Indeed, the set $\{\tilde{B}_{k}\mid k\in[n]\}$ is in distribution equal $n$ independent standard Brownian motions, since $\operatorname{\mathbb{E}}\left[\mathrm{d}\tilde{B}_{k}\right]=0$ and for $k,l\in[n]$
$\operatorname{\mathbb{E}}\left[\mathrm{d}\tilde{B}_{k}\mathrm{d}\tilde{B}_{l}\right]$ $=\operatorname{\mathbb{E}}\left[\frac{1}{2}\sum_{ij}\sum_{st}\sqrt{1+\delta_{ij}}\sqrt{1+\delta_{st}}(v_{k})_{i}(v_{k})_{j}(v_{l})_{s}(v_{l})_{t}\mathrm{d}B_{ij}\mathrm{d}B_{st}\right]$
where we realize that the product of the differentials is 0 except for $(i,j)=(s,t)$ and $(i,j)=(t,s)$
$=\operatorname{\mathbb{E}}\left[v_{k}^{T}v_{l}v_{k}^{T}v_{l}\right]\mathrm{d}t$
$=\delta_{kl}\mathrm{d}t.$ (44)
We can thus conclude that by eq. 20 we have,
$\mathrm{d}\lambda_{k}=\left(-\beta\lambda_{k}+\alpha\sum_{l\in[n]\setminus\{k\}}\frac{1}{\lambda_{k}-\lambda_{l}}\right)\mathrm{d}t+\sqrt{2\alpha}\mathrm{d}W_{k}$ (45)
where $\{W_{k}\mid k\in[n]\}$ are $n$ independent Brownian motions.
Invariant Distribution.
We now show that
$p_{\mathrm{inv}}(\lambda)=\frac{1}{Z}\exp(-U(\lambda))\quad\quad\text{ for }\quad\quad U(\lambda)=\frac{\beta}{2\alpha}\sum_{k}\lambda_{k}^{2}-\sum_{k<\ell}\ln|\lambda_{k}-\lambda_{\ell}|,$ (46)
for $\lambda\in C_{n}$ and $Z$ a normalizing constant so that $p_{\mathrm{inv}}$ is a probability measure, is the invariant distribution. We use the Fokker-Plank equation. Indeed, recall that if we have an SDE
$\mathrm{d}\lambda_{t}=f(\lambda_{t})\mathrm{d}t+L(\lambda_{t})\mathrm{d}B_{t}$
in dimension $n$, where $f(\lambda)=(f_{1}(\lambda),\ldots,f_{n}(\lambda))$ is a $C^{2}$-function, $L(\lambda)$ is matrix-valued $C^{2}$-function and $B_{t}$ is a $n$-dimensional Brownian motion, then an invariant distribution $p(x)$ satisfies
$\sum_{i=1}^{n}\frac{\partial}{\partial\lambda_{i}}[f_{i}(\lambda)p(\lambda)]=\frac{1}{2}\sum_{i,j=1}^{n}\frac{\partial^{2}}{\partial\lambda_{i}\partial\lambda_{j}}[L(\lambda)L(\lambda)^{T}]_{ij}p(\lambda).$ (47)
In our case, for the $\lambda(\alpha,\beta)$-SDE we have that $L(\lambda,t)=\sqrt{2\alpha}1_{n}$ and therefore $L(\lambda,t)L(\lambda,t)^{T}]_{ij}=2\alpha\delta_{ij}$. So the right hand side of (47) equals
$\alpha\sum_{i=1}^{n}\frac{\partial^{2}}{(\partial\lambda_{i})^{2}}p(\lambda).$
Assume for now that $\lambda=(\lambda_{1},\ldots,\lambda_{n})$ satisfies $\lambda_{1}>\lambda_{2}>\ldots>\lambda_{n}$ and
$U(\lambda)=c\sum_{i}\lambda_{i}^{2}-\sum_{i<j}\ln(\lambda_{i}-\lambda_{j})$
for some constant $c>0$ to be determined. We first calculate for a fixed $i$,
$\frac{\partial}{\partial\lambda_{i}}p(\lambda)$ $=-\frac{1}{Z}\exp(-U(\lambda))\frac{\partial}{\partial\lambda_{i}}U(\lambda)$
$=-\frac{1}{Z}\exp(-U(\lambda))\left(2c\lambda_{i}-\sum_{i<j}\frac{1}{\lambda_{i}-\lambda_{j}}+\sum_{j<i}\frac{1}{\lambda_{j}-\lambda_{i}}\right)$
$=-\frac{1}{Z}\exp(-U(\lambda))\left(2c\lambda_{i}-\sum_{j\neq i}\frac{1}{\lambda_{i}-\lambda_{j}}\right).$
Therefore,
$\frac{\partial^{2}}{(\partial\lambda_{i})^{2}}p(\lambda)$ $=\frac{1}{Z}\exp(-U(\lambda))\left(2c\lambda_{i}-\sum_{j\neq i}\frac{1}{\lambda_{i}-\lambda_{j}}\right)^{2}-\frac{1}{Z}\exp(-U(\lambda))\left(2c+\sum_{j\neq i}\frac{1}{(\lambda_{i}-\lambda_{j})^{2}}\right)$
$=\left(\sum_{j\neq i}\frac{1}{\lambda_{i}-\lambda_{j}}-2c\lambda_{i}\right)\frac{\partial}{\partial\lambda_{i}}p(\lambda)-p(\lambda)\left(2c+\sum_{j\neq i}\frac{1}{(\lambda_{i}-\lambda_{j})^{2}}\right)$
and so the right hand side of (47) is equal to
$\alpha\sum_{i=1}^{n}\frac{\partial^{2}}{(\partial\lambda_{i})^{2}}p(\lambda)=\sum_{i=1}^{n}\left(\alpha\sum_{j\neq i}\frac{1}{\lambda_{i}-\lambda_{j}}-2\alpha c\lambda_{i}\right)\frac{\partial}{\partial\lambda_{i}}p(\lambda)-\sum_{i=1}^{n}p(\lambda)\left(2\alpha c+\alpha\sum_{j\neq i}\frac{1}{(\lambda_{i}-\lambda_{j})^{2}}\right)$
Now the left hand side of (47) is equal to
$\sum_{i=1}^{n}\frac{\partial}{\partial\lambda_{i}}\left[\left(\alpha\sum_{i\neq j}\frac{1}{\lambda_{i}-\lambda_{j}}-\beta\lambda_{i}\right)p(\lambda)\right]$
$=$ $\sum_{i=1}^{n}\left(\alpha\sum_{i\neq j}\frac{1}{\lambda_{i}-\lambda_{j}}-\beta\lambda_{i}\right)\frac{\partial}{\partial\lambda_{i}}p(\lambda)+\sum_{i=1}^{n}\left(-\alpha\sum_{i\neq j}\frac{1}{(\lambda_{i}-\lambda_{j})^{2}}-\beta\right)p(\lambda)$
So it follows that in (47) the left-hand side is equal to the right-hand side if and only if $\beta=2\alpha c$ or equivalently $c=\beta/2\alpha$, concluding the proof. ∎
Figure 4: Plot of the invariant density of Dyson-BM for $d=2$ and $\alpha=\beta=1$
Appendix C Inferring the Eigenvector Dynamics
Theorem 3.2 (restated).
Denote by $(v_{1}(t),\ldots,v_{n}(t))$ the orthonormal eigenvectors associated to the eigenvalues of Theorem 3.1. Assuming that the initial matrix $M(0)$ has simple spectrum, $v_{k}(t)$ satisfies for $k\in[n]$ the stochastic differential equation
$\mathrm{d}v_{k}(t)=-\frac{\alpha}{2}\sum_{\ell\neq k}\frac{1}{(\lambda_{k}(t)-\lambda_{\ell}(t))^{2}}v_{k}(t)\mathrm{d}t+\sqrt{\alpha}\sum_{\ell\neq k}\frac{1}{\lambda_{k}(t)-\lambda_{\ell}(t)}v_{\ell}(t)\mathrm{d}w_{\ell k}(t)$ (Eigenvector-SDE)
for $\{w_{ij:i\neq j}\}$ standard Brownian motions also independent of the eigenvalue trajectories, with $w_{ji}=w_{ij}$.
Proof.
Analogously to the proof of Theorem 3.1, we may view the $v_{k}$ for $k\in[n]$ as a function of the matrix components $M_{ij}$ for $(i,j)\in\mathcal{I}\coloneqq\{(i,j):1\leq i\leq j\leq n\}$. We thus have by Itô’s lemma
$\mathrm{d}v_{k}$ $=\sum_{\eta\in\mathcal{I}}\frac{\partial v_{k}}{\partial M_{\eta}}\mathrm{d}M_{\eta}+\frac{1}{2}\sum_{\eta,\xi\in\mathcal{I}}\frac{\partial^{2}v_{k}}{\partial M_{\eta}\partial M_{\xi}}\mathrm{d}M_{\eta}\mathrm{d}M_{\xi}$
$=\sum_{\eta\in\mathcal{I}}\left(-\beta M_{\eta}\frac{\partial v_{k}}{\partial M_{\eta}}+\frac{1}{2}D_{\eta}^{2}\frac{\partial^{2}v_{k}}{(\partial M_{\eta})^{2}}\right)dt+D_{\eta}\frac{\partial v_{k}}{\partial M_{\eta}}\mathrm{d}B_{\eta}.$ (48)
For the first summand of eq. 48, we observe, using eq. 36, that
$\sum_{(i,j)\in\mathcal{I}}M_{ij}\frac{\partial v_{k}}{\partial M_{ij}}\mathrm{d}t=\sum_{l\neq k}\sum_{(i,j)\in\mathcal{I}}v_{l}^{T}\frac{\partial v_{k}}{\partial M_{ij}}v_{l}M_{ij}\mathrm{d}t$ (49)
and further use eq. 30
[…]
$=$ $\sum_{l\neq k}\frac{\lambda_{k}}{\lambda_{k}-\lambda_{l}}(v_{l}^{T}v_{k})v_{l}\mathrm{d}t.$ (50)
Since $l\neq k$ and the $\{v_{l}\mid l\in[n]\}$ are orthogonal, we get
$=$ $0.$ (51)
For the second summand of eq. 48, we observe
$\frac{\partial^{2}v_{k}}{(\partial M_{\eta})^{2}}=\frac{\partial}{\partial M_{\eta}}\frac{\partial v_{k}}{\partial M_{\eta}}$ (52)
and using Equation 30 and Equation 36 we get
$=$ $\sum_{m\neq k}\frac{\partial}{\partial M_{\eta}}\left[\frac{1}{\lambda_{k}-\lambda_{m}}v_{m}^{T}\frac{\partial M}{\partial M_{\eta}}v_{k}v_{m}\right]$ (53)
by noting $\frac{\partial}{\partial M_{\eta}}\frac{\partial M}{\partial M_{\eta}}=0$, we apply the chain rule to obtain
$=$ $\sum_{m\neq k}\Bigg\{-\underbrace{\left(\frac{1}{\lambda_{k}-\lambda_{m}}\right)^{2}\left(\frac{\partial\lambda_{k}}{\partial M_{\eta}}-\frac{\partial\lambda_{m}}{\partial M_{\eta}}\right)v_{m}^{T}\frac{\partial M}{\partial M_{\eta}}v_{k}v_{m}}_{(i)}+\underbrace{\frac{1}{\lambda_{k}-\lambda_{m}}\frac{\partial v_{m}^{T}}{\partial M_{\eta}}\frac{\partial M}{\partial M_{\eta}}v_{k}v_{m}}_{(ii)}$ (54)
$\quad+\underbrace{\frac{1}{\lambda_{k}-\lambda_{m}}v_{m}^{T}\frac{\partial M}{\partial M_{\eta}}\frac{\partial v_{k}}{\partial M_{\eta}}v_{m}}_{(iii)}+\underbrace{\frac{1}{\lambda_{k}-\lambda_{m}}v_{m}^{T}\frac{\partial M}{\partial M_{\eta}}v_{k}\frac{\partial{v_{m}}}{\partial M_{\eta}}}_{(iv)}\Bigg\}.$ (55)
We now analyze each term.
Term (i)
We recall Equation 32
$\frac{\partial\lambda_{k}}{\partial M_{ij}}$ $=(v_{k})_{i}(v_{k})_{j}(2-\delta_{ij}),$
as well as that the partial derivative $\frac{\partial M}{\partial M_{ij}}$ is 0 except at $i,j$ and $j,i$, where it is $1$ (using symmetry).
Pulling the summation over $\eta$ and all $\eta$-dependent terms in, we have
$\sum_{\eta\in\mathcal{I}}(1+\delta_{\eta})\left(\frac{\partial\lambda_{k}}{\partial M_{\eta}}-\frac{\partial\lambda_{m}}{\partial M_{\eta}}\right)v_{m}^{T}\frac{\partial M}{\partial M_{\eta}}v_{k}$
$=$ $\sum_{ij\in\mathcal{I}}(1+\delta_{ij})\Bigg\{\left[(v_{k})_{i}(v_{k})_{j}(2-\delta_{ij})-(v_{m})_{i}(v_{m})_{j}(2-\delta_{ij})\right]$
$\qquad\left[(v_{m})_{i}(v_{k})_{j}+(v_{m})_{j}(v_{k})_{i}(1-\delta_{ij})\right]\Bigg\}$
$=$ $2\left[v_{k}v_{k}^{T}v_{k}^{T}v_{m}-v_{m}^{T}v_{m}v_{m}^{T}v_{k}\right]$
$=$ $2(\delta_{km}-\delta_{km})$
$=$ $0.$
Thus, the overall contribution of term $(i)$ is 0.
Term (ii)
For term (ii) we have by Equation 30 and Equation 36
$(ii)=$ $\frac{1}{\lambda_{k}-\lambda_{m}}\sum_{s\neq m}\frac{1}{\lambda_{m}-\lambda_{s}}v_{m}^{T}\frac{\partial M}{\partial M_{\eta}}v_{s}v_{s}^{T}\frac{\partial M}{\partial M_{\eta}}v_{k}v_{m}.$ (56)
Using
$v_{m}^{T}\frac{\partial M}{\partial M_{ij}}v_{s}=(v_{m})_{i}(v_{s})_{j}+(v_{m})_{j}(v_{s})_{i}(1-\delta_{ij}),$ (57)
we note that
$\sum_{\eta\in\mathcal{I}}D_{\eta}^{2}v_{m}^{T}\frac{\partial M}{\partial M_{\eta}}v_{s}v_{s}^{T}\frac{\partial M}{\partial M_{\eta}}v_{k}$
$=$ $\alpha\sum_{i,j\in\mathcal{I}}(1+\delta_{ij})\left[(v_{m})_{i}(v_{s})_{j}+(v_{m})_{j}(v_{s})_{i}(1-\delta_{ij})\right]\left[(v_{k})_{i}(v_{s})_{j}+(v_{k})_{j}(v_{s})_{i}(1-\delta_{ij})\right]$
$=$ $\alpha\left\{\left[\sum_{i}(v_{m})_{i}(v_{k})i\right]\left[\sum_{i}(v_{s})_{i}^{2}\right]+v_{m}^{T}v_{s}v_{s}^{T}v_{k}\right\}$
$=$ $\alpha\delta_{km}+\alpha\delta_{ms}\delta_{sk}$ (58)
$=$ $0.$
Thus, the contribution of term $(ii)$ vanishes.
Term (iii)
Analogous to term $(ii)$.
Term (iv)
By using Equation 30 and Equation 36, we get
$(iv)=$ $\frac{1}{\lambda_{k}-\lambda_{m}}v_{m}^{T}\frac{\partial M}{\partial M_{\eta}}v_{k}\frac{\partial{v_{m}}}{\partial M_{\eta}}$
$=$ $\frac{1}{\lambda_{k}-\lambda_{m}}v_{m}^{T}\frac{\partial M}{\partial M_{\eta}}v_{k}\sum_{p\neq m}\frac{1}{\lambda_{m}-\lambda_{p}}v_{p}^{T}\frac{\partial M}{\partial M_{\eta}}v_{m}v_{p}$
$=$ $\frac{1}{\lambda_{k}-\lambda_{m}}\sum_{p\neq m}\frac{1}{\lambda_{m}-\lambda_{p}}\Big(v_{m}^{T}\frac{\partial M}{\partial M_{\eta}}v_{k}\Big)\Big(v_{p}^{T}\frac{\partial M}{\partial M_{\eta}}v_{m}\Big)v_{p}.$
By pulling in the $\eta$-dependent terms from Equation 48 which affect term $(iv)$, and then by subsequently pulling the sum over $\eta$ in, we get
$\sum_{\eta\in\mathcal{I}}\frac{1}{2}D_{\eta}^{2}v_{m}^{T}\frac{\partial M}{\partial M_{\eta}}v_{k}v_{p}^{T}\frac{\partial M}{\partial M_{\eta}}v_{m}$ (59)
which, using eq. 58, becomes
$=$ $\frac{\alpha}{2}\left(\delta_{kp}+\delta_{km}\delta_{mp}\right)$
$=$ $\frac{\alpha}{2}\delta_{kp}.$
Moreover, using that $p\neq m$, we can conclude that the total contribution of term $(iv)$ is
$\sum_{\eta\in\mathcal{I}}\frac{1}{2}D_{\eta}^{2}\sum_{m\neq k}(iv)$
$=$ $-\frac{\alpha}{2}\sum_{m\neq k}\frac{1}{(\lambda_{k}-\lambda_{m})^{2}}v_{k}.$
Before proceeding, we prove the following lemma.
Lemma C.1.
Let $v^{(k)}$ be a set of orthonormal vectors and $B_{ij}$ a set of independent standard Brownian motions with $B_{ij}=B_{ji}$, for $i,j,k\in[n]$. We have that
$\tilde{B}_{lk}:=(v^{(l)})^{T}\begin{pmatrix}\sqrt{2}B_{11}&B_{12}&B_{13}&\ldots&B_{1n}\\ B_{21}&\sqrt{2}B_{22}&B_{23}&\ldots&B_{2n}\\ \vdots&&\vdots&&\vdots\\ B_{n1}&B_{n2}&B_{n3}&\ldots&\sqrt{2}B_{nn}\\ \end{pmatrix}v^{(k)}$
is a Brownian motion with variance $1+\delta_{lk}$, that is $\tilde{B}_{lk}\stackrel{{\scriptstyle\rm distr}}{{=}}\sqrt{1+\delta_{lk}}B$ for $B$ a standard Brownian motion. $\tilde{B}_{lk}$ is independent of $\tilde{B}_{ab}$ for $(l,k)\neq(a,b)$ and $(l,k)\neq(b,a)$.
Proof.
We have for any $a,b\in[n]$ that
$\mathrm{d}\tilde{B}_{ab}=\sum_{i,j}(v^{(a)})_{i}(v^{(b)})_{j}\sqrt{1+\delta_{ij}}\mathrm{d}B_{ij}.$
Thus, $\operatorname{\mathbb{E}}\left[\mathrm{d}\tilde{B}_{ab}\right]=0$. We further have for $c,d\in[n]$:
$\operatorname{\mathbb{E}}\left[\mathrm{d}\tilde{B}_{ab}\mathrm{d}\tilde{B}_{cd}\right]$ $=\sum_{ijkl}(v^{(a)})_{i}(v^{(b)})_{j}(v^{(c)})_{k}(v^{(d)})_{l}\sqrt{1+\delta_{ij}}\sqrt{1+\delta_{kl}}\operatorname{\mathbb{E}}\left[\mathrm{d}B_{ij}\mathrm{d}B_{kl}\right]$
$=\sum_{ijkl}(v^{(a)})_{i}(v^{(b)})_{j}(v^{(c)})_{k}(v^{(d)})_{l}\sqrt{1+\delta_{ij}}\sqrt{1+\delta_{kl}}\mathds{1}_{(i,j)=(k,l)\lor(i,j)=(l,k)}\mathrm{d}t$
$=\sum_{ij}(v^{(a)})_{i}(v^{(b)})_{j}\left\{\mathds{1}_{i\neq j}\left[(v^{(c)})_{i}(v^{(d)})_{j}+(v^{(c)})_{j}(v^{(d)})_{i}\right]+\mathds{1}_{i=j}(v^{(c)})_{i}(v^{(d)})_{i}\right\}(1+\delta_{ij})\mathrm{d}t$
$=\sum_{ij}(v^{(a)})_{i}(v^{(b)})_{j}\left[(v^{(c)})_{i}(v^{(d)})_{j}+(v^{(c)})_{j}(v^{(d)})_{i}\right]\mathrm{d}t$
$=\sum_{ij}(v^{(a)})_{i}(v^{(b)})_{j}(v^{(c)})_{i}(v^{(d)})_{j}\mathrm{d}t+\sum_{ij}(v^{(a)})_{i}(v^{(b)})_{j}(v^{(c)})_{j}(v^{(d)})_{i}\mathrm{d}t$
$=(v^{(a)})^{T}v^{(c)}(v^{(b)})^{T}v^{(d)}\mathrm{d}t+(v^{(a)})^{T}v^{(d)}(v^{(b)})^{T}v^{(c)}\mathrm{d}t$
$=\mathds{1}_{(a,b)=(c,d)}\mathrm{d}t+\mathds{1}_{(a,b)=(d,c)}\mathrm{d}t$ (60)
Thus, in particular the process $\tilde{B}_{ab}$ has variance $t$ for $a\neq b$ and variance $2t$ for $a=b$.
Since the linear combination of independent Brownian Motions is jointly normal, we see from the Covariance property in eq. 60 that $\tilde{B}_{lk}$ is independent of $\tilde{B}_{ab}$ for $(l,k)\neq(a,b)$ and $(l,k)\neq(b,a)$. ∎
We can now turn to the third summand of eq. 48.
We have
$\sum_{\eta\in\mathcal{I}}D_{\eta}\frac{\partial v_{k}}{\partial M_{\eta}}\mathrm{d}B_{\eta}$
$=$ $\sqrt{\alpha}\sum_{l\neq k}\frac{1}{\lambda_{k}-\lambda_{l}}v_{l}^{T}\sum_{\eta\in\mathcal{I}}\frac{\partial M}{\partial M_{\eta}}v_{k}\sqrt{1+\delta_{\eta}}\mathrm{d}B_{\eta}v_{l}$
$=$ $\sqrt{\alpha}\sum_{l\neq k}\frac{1}{\lambda_{k}-\lambda_{l}}v_{l}^{T}\begin{pmatrix}\sqrt{2}\mathrm{d}B_{11}&\mathrm{d}B_{12}&\mathrm{d}B_{13}&\ldots&\mathrm{d}B_{1n}\\ \mathrm{d}B_{21}&\sqrt{2}\mathrm{d}B_{22}&\mathrm{d}B_{23}&\ldots&\mathrm{d}B_{2n}\\ \vdots&&\vdots&&\vdots\\ \mathrm{d}B_{n1}&\mathrm{d}B_{n2}&\mathrm{d}B_{n3}&\ldots&\sqrt{2}\mathrm{d}B_{nn}\\ \end{pmatrix}v_{k}v_{l}$
$=$ $\sqrt{\alpha}\sum_{l\neq k}\frac{1}{\lambda_{k}-\lambda_{l}}v_{l}\mathrm{d}\tilde{B}_{lk}$ (61)
where we used Lemma C.1 in the last step so that $\mathrm{d}\tilde{B}_{lk}$ are symmetric Brownian motions with variance $1+\delta_{lk}$.
Having established all the terms in the SDE in Theorem 3.2, we check that this dynamics of the eigenvectors gives rise to normalized vectors, assuming that the initial vectors $(v_{1}(0),\dots,v_{n}(0))$ are normalized. To this end, it suffices to note that $v_{k}(t+\mathrm{d}t)=v_{k}(t)+dv_{k}(t)$ is normalized given that $v_{k}(t)$ is normalized, that is $(v_{k}(t))^{2}=1$ for any $t$. By continuity of $t\mapsto v_{k}(t)$, it suffices to show that $v_{k}(\mathrm{d}t)=v_{k}(0)+dv_{k}(0)$ remains normalized. We use the SDE from Theorem 3.2 to compute that up terms of the order of $\mathrm{d}t^{3/2}$ or higher we have (omitting “(0)” in the notation)
$(v_{k}(0)+\mathrm{d}v_{k}(0))^{2}=1+2v_{k}^{T}\mathrm{d}v_{k}+(\mathrm{d}v_{k})^{2}$
$=$ $1+2\left(\frac{-\alpha}{2}\right)\sum_{l\neq k}\frac{\mathrm{d}t}{(\lambda_{k}-\lambda_{l})^{2}}+2\sqrt{\alpha}\sum_{l\neq k}\frac{1}{\lambda_{l}-\lambda_{k}}\underbrace{v_{k}^{T}v_{l}}_{=0}\mathrm{d}w_{lk}$
$+\alpha\sum_{l\neq k}\sum_{j\neq k}\frac{\mathrm{d}w_{lk}\mathrm{d}w_{jk}}{(\lambda_{l}-\lambda_{k})(\lambda_{j}-\lambda_{k})}+\mathcal{O}(\mathrm{d}t^{3/2})$
$=$ $1-\alpha\sum_{l\neq k}\frac{\mathrm{d}t}{(\lambda_{k}-\lambda_{l})^{2}}+\alpha\sum_{l\neq k}\sum_{j\neq k}\frac{\delta_{jl}\mathrm{d}t}{(\lambda_{k}-\lambda_{l})(\lambda_{k}-\lambda_{j})}+\mathcal{O}(\mathrm{d}t^{3/2})$
$=$ $1+\mathcal{O}(\mathrm{d}t^{3/2}).$ (62)
Since contributions of $\mathcal{O}(\mathrm{d}t^{3/2})$ do not contribute to the dynamics for $\mathrm{d}t\to 0$, this implies that the $v_{k}(\mathrm{d}t)$ remain normalized. The argument may be iterated, i.e., $(v_{k}(2\mathrm{d}t))^{2}=(v_{k}(\mathrm{d}t)+\mathrm{d}v_{k}(\mathrm{d}t))^{2}=1+\mathcal{O}(\mathrm{d}t^{3/2})$ and so on, and thus the $v_{k}(t)$ remain normalized for all $t>0$.
Finally, as an instructive consistency check, we also show that the eigenvectors remain orthogonal to each other (as they must, since $M$ is symmetric for all $t\geq 0$). This can be checked from the SDE in an analogous manner using again the continuity of $t\mapsto v_{k}(t)$. Note that for $k\neq l$ and given $v_{k}^{T}(0)v_{l}(0)=0$, it suffices to show that $v_{k}^{T}(\mathrm{d}t)v_{l}(\mathrm{d}t)=0$. Indeed, we have (omitting “(0)” in the notation)
$(v_{k}(0)+\mathrm{d}v_{k}(0))^{T}(v_{l}(0)+\mathrm{d}v_{l}(0))=0+v_{k}^{T}dv_{l}+v_{l}^{T}dv_{k}+\mathrm{d}v_{k}^{T}dv_{l}$
$=$ $\sqrt{\alpha}\sum_{i\neq l}\frac{1}{\lambda_{l}-\lambda_{i}}v_{k}^{T}v_{i}\mathrm{d}w_{il}+\sqrt{\alpha}\sum_{j\neq k}\frac{1}{\lambda_{k}-\lambda_{j}}v_{l}^{T}v_{j}\mathrm{d}w_{jk}$
$+\alpha\sum_{j\neq k}\sum_{i\neq l}\frac{v_{j}^{T}v_{i}\mathrm{d}w_{il}\mathrm{d}w_{jk}}{(\lambda_{k}-\lambda_{j})(\lambda_{l}-\lambda_{i})}+\mathcal{O}(\mathrm{d}t^{3/2})$
$=$ $\frac{\sqrt{\alpha}}{\lambda_{l}-\lambda_{k}}\underbrace{(\mathrm{d}w_{kl}-\mathrm{d}w_{lk})}_{=0}+\alpha\sum_{j\neq k}\sum_{i\neq l}\frac{\delta_{ij}\delta_{ik}\delta_{jl}\mathrm{d}t}{(\lambda_{k}-\lambda_{j})(\lambda_{l}-\lambda_{i})}+\mathcal{O}(\mathrm{d}t^{3/2})$
$=$ $0+\mathcal{O}(\mathrm{d}t^{3/2}).$ (63)
The argument may be iterated and thus the claim hold for all $t>0$.
Appendix D Time reversal: Existence and Uniqueness
The time-reversal of the Dyson-BM in the sense of Ref. Anderson (1982) is given in Eq. (5). Since the drift coefficient in the Dyson-BM is not locally Lipschitz continuous, the existence and uniqueness of strong solutions to Dyson-BM and Eq. (5), and the applicability of Ref. Anderson (1982) are not obvious. While the existence and uniqueness of a strong solution to the Dyson-BM is well established (see, e.g., Lemma 4.3.3 in Ref. Anderson et al. (2009)), we here sketch how to ensure the other two points.
To ensure existence and uniqueness of a strong solution to Eq. (5), repeat the arguments in Lemma 4.3.3 in Ref. Anderson et al. (2009), where the divergent $x^{-1}$-term in the drift is replaced by the locally Lipschitz continuous approximation $\phi(x)=x^{-1}$ for $|x|\geq R^{-1}$ and $\phi(x)=xR^{2}$ otherwise. For any $R>0$, the desired statements follow from the local Lipschitz continuity of the drift, and details on the limit $R\to 0$ can be found in Ref. Anderson et al. (2009). The only difference to the forward motion is the additional drift term containing the score function. However, since the dynamics is almost surely contained in the interior of the Weyl chamber (Katori & Tanemura, 2003), the propagator in the score-contribution is, as usual, dominated by white noise as $\mathrm{d}t\to 0$. Therefore, this term will not cause complications as $R\to 0$ and the arguments from Ref. Anderson et al. (2009) imply uniqueness and existence of a strong solution to Eq. (5).
Using the same regularization of the divergent term, for any $R>0$ the statements of Ref. Anderson (1982) are directly applicable. Since this regularization is piecewise continuous, we can take the limit $R\to 0$ under time reversal. Since we just established the existence of the solution to Eq. (5), the limit $R\to 0$ converges. Furthermore, the arguments are expected to generalize (regularize $\phi(x)=x^{-2}$) to the eigenvector equations and their time reversal.
Appendix E Time rescaling
We rescale time for Dyson-BM. Let $T(s)$ be a continuous, differentiable rescaling of time, monotonically increasing, with $T(0)=0$ (for convenience).
Let $\gamma_{k}(t):=\lambda(T(t))$. With this definition, the Ito SDE eq. Dyson-BM corresponds to the Ito integral
$\gamma_{k}(t):=$ $\lambda_{k}(T(t))=\lambda_{k}(0)+\int_{0}^{T(t)}\left(\alpha\sum_{i\neq k}\frac{1}{\lambda_{k}(s)-\lambda_{i}(s)}-\beta\lambda_{k}(s)\right)\mathrm{d}s+\int_{0}^{T(t)}\sqrt{2\alpha}\mathrm{d}B_{k}(s)$
$=$ $\int_{0}^{t}\left(\alpha\sum_{i\neq k}\frac{1}{\lambda_{k}(T(s))-\lambda_{i}(T(s))}-\beta\lambda_{k}(T(s))\right)T^{\prime}(s)\mathrm{d}s+\int_{0}^{t}\sqrt{2\alpha T^{\prime}(s)}\mathrm{d}B_{k}(s)$ (using Thm. 8.5.7 in Øksendal (2003))
$=$ $\int_{0}^{t}\left(\alpha\sum_{i\neq k}\frac{1}{\gamma_{k}(s)-\gamma_{i}(s)}-\beta\gamma_{k}(s)\right)T^{\prime}(s)\mathrm{d}s+\int_{0}^{t}\sqrt{2\alpha T^{\prime}(s)}\mathrm{d}B_{k}(s)$ (64)
which we can write in SDE notation
$\mathrm{d}\gamma_{k}(t)=\left(\alpha\sum_{i\neq k}\frac{1}{\gamma_{k}(t)-\gamma_{i}(t)}-\beta\gamma_{k}(t)\right)T^{\prime}(t)\mathrm{d}t+\sqrt{2\alpha T^{\prime}(t)}\mathrm{d}B_{k}(t).$ (65)
By using $T(t):=\frac{1}{\alpha}t$, we obtain
$\mathrm{d}\gamma_{k}(t)=\left(\sum_{i\neq k}\frac{1}{\gamma_{k}(t)-\gamma_{i}(t)}-\frac{\beta}{\alpha}\gamma_{k}(t)\right)\mathrm{d}t+\sqrt{2}\mathrm{d}B_{k}(t),$ (66)
so that we can summarize the two parameters to $\eta:=\frac{\beta}{\alpha}$:
$\mathrm{d}\gamma_{k}(t)=\left(\sum_{i\neq k}\frac{1}{\gamma_{k}(t)-\gamma_{i}(t)}-\eta\gamma_{k}(t)\right)\mathrm{d}t+\sqrt{2}\mathrm{d}B_{k}(t)$ (67)
which we call $\gamma(\eta)$-SDE”.
Dyson’s conjecture says that the $\lambda(\frac{1}{N},\frac{1}{2})$-SDE converges to global equilibrium in time $\Theta(1)$ (see Yang (2022)). Running $\lambda(\frac{1}{N},\frac{1}{2})$ until time 1 is the same as running the $\gamma(\frac{N}{2})$-SDE until time $T(1)=\frac{1}{N}$.
Appendix F Stepsize Controller
Dyson Brownian Motion almost surely never crosses the singularities. Hence, conditioning on non-crossing corresponds to conditioning on a probability $1$ event, which does not change the dynamics. Given the noise, we can thus calculate the maximal step size, beyond which we would cross the singularity. This is a very useful upper bound, which we employ in practice to get the numerical scheme working. It has two effects: (1) close to the boundary of the Weyl Chamber, it avoids numerically stepping over the singularities and (2) far from the boundary, it allows for larger step size, increasing efficiency.
F.1 Forward in Time
Algorithm 2 Forward stepsize controller for Dyson SDE
1:Input: position $\lambda\in\mathbb{R}^{n}$, time $t\in\mathbb{R}^{+}$, independent normal samples $u\sim\mathcal{N}^{n}$.
2:for $k\in[n-1]$ do
3:   $\delta t_{k}\leftarrow$ maximal step size based on $\lambda_{k+1}-\lambda_{k}$ and samples $u_{k}$, $u_{k+1}$ as described in Section F.1
4:end for
5:Output: stepsize $\min_{i}\delta t_{i}$.
Figure 5: Forward step size controller which exploits that non-crossing of paths happens with probability 11. The exact calculations are carried out in Section F.1.
The difference between components $k$ and $k+1$ of $\lambda$ after a first-order discretization step of size $\mathrm{d}t$ reads for any time $t\in\mathbb{R}^{+}$
$\Delta_{k}(t+\mathrm{d}t)$ $\coloneqq\lambda_{k}(t+\mathrm{d}t)-\lambda_{k+1}(t+\mathrm{d}t)$ (68)
$=\lambda_{k}(t)-\lambda_{k+1}(t)+\mathrm{d}tf_{k}(t)+\sqrt{\mathrm{d}t}g_{k}$ (69)
with functions $f_{k}(t)\coloneqq\alpha\left(\sum_{i\neq k}\frac{1}{\lambda_{k}(t)-\lambda_{i}(t)}-\sum_{i\neq k+1}\frac{1}{\lambda_{k+1}(t)-\lambda_{i}(t)}\right)-\beta\left(\lambda_{k}(t)-\lambda_{k+1}(t)\right)$, $g_{k}\coloneqq\sqrt{2\alpha}\left(X_{k}-X_{k+1}\right)$ where the $X_{j}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,1)$ are the independent standard gaussian random variables taken in the update step of the numerical SDE scheme. Note that $\Delta_{k}$ is a random variable, but if the step size is sufficiently small we must have almost surely
$\Delta_{k}(t+\mathrm{d}t)$ $>0.$
To find the maximal step size, we observe that the equation above is a quadratic function in the substituted $\tau:=\sqrt{\mathrm{d}t}$ yielding the inequality that
$\tau^{2}+\frac{g_{k}}{f_{k}(t)}\tau+\frac{\Delta_{k}(t)}{f_{k}(t)}\qquad\text{is }\qquad\begin{cases}>0&\text{if $f_{k}(t)>0$},\\ <0&\text{if $f_{k}(t)<0$.}\end{cases}$ (70)
We treat first the case $f_{k}(t)>0$. The inequality is equivalent to
$\left(\tau+\frac{g_{k}}{2f_{k}(t)}\right)^{2}>\left(\frac{g_{k}}{2f_{k}(t)}\right)^{2}-\frac{\Delta_{k}(t)}{f_{k}(t)},$ (71)
where we know that $\frac{\Delta_{k}(t)}{f_{k}(t)}>0$ must hold. Hence, if $g_{k}>0$, any stepsize $\mathrm{d}t>0$ works. Otherwise, if $g_{k}<0$, the quadratic formula shows immediately that both roots will be in the $\tau>0$ regime. Since the parabola is convex in $\tau$, for inequality (71) to be fulfilled, we take $\tau$ in the range from 0 to the smallest root. For $\mathrm{d}t$, that means
$\mathrm{d}t\in\left(0,\frac{1}{4}\left(-\frac{g_{k}}{f_{k}(t)}-\sqrt{\frac{g_{k}^{2}}{f_{k}(t)^{2}}-4\frac{\Delta_{k}(t)}{f_{k}(t)}}\right)^{2}\right).$ (72)
If $f_{k}(t)<0$, we have
$\left(\tau+\frac{g_{k}}{2f_{k}(t)}\right)^{2}<\left(\frac{g_{k}}{2f_{k}(t)}\right)^{2}-\frac{\Delta_{k}(t)}{f_{k}(t)}.$ (73)
Since $\frac{\Delta_{k}(t)}{f_{k}(t)}<0$, we know that the inequality is satisfied at $\tau=0$, and that the largest root $\tau_{2}$ will be positive:
$\tau_{1,2}=\frac{1}{2}\left(-\frac{g_{k}}{f_{k}(t)}\mp\sqrt{\frac{g_{k}^{2}}{f_{k}(t)^{2}}-4\frac{\Delta_{k}(t)}{f_{k}(t)}}\right).$
Thus, any stepsize in the following interval is valid
$\mathrm{d}t\in\left(0,\frac{1}{4}\left(-\frac{g_{k}}{f_{k}(t)}+\sqrt{\frac{g_{k}^{2}}{f_{k}(t)^{2}}-4\frac{\Delta_{k}(t)}{f_{k}(t)}}\right)^{2}\right)$
Note that $f_{k}(t)\neq 0$ and $g_{k}\neq 0$ almost surely.
F.2 Backward in Time
Backwards in time, we carry out the analogous computation for the more involved backwards SDE in eq. 5. In essence, this boils again down to solving a quadratic equation and considering all edge cases.
Appendix G Shooting mechanism
In the forward dynamics, if a certain step size would lead to the probability-0 event of leaving the Weyl-Chamber, we know that the source of the error is the finite-time-step discrete approximation, so that decreasing the step size will always provide a fix. In the backward dynamics, however, this is not guaranteed: As described in Section 3.2, a possible reason for this probability-0 event in the backwards dynamics is that the score $s_{\theta}$ might be not perfectly learned: $s_{\theta}(\lambda,t)\neq s(\lambda,t)$ for some $\lambda,t$. To avoid this probability-0 event, we use the following “shooting mechanism”: If we were to leave the Weyl-Chamber, we replace the learned score with the analytically known score in the invariant state, eliminating the use of the neural network at that point and leading effectively to a repulsion with the negative forward drift. This mechanism is not expected to change the dynamics in any unfavorable way since it is only applied very rarely, and only in cases where the actual learned dynamics is a much worse approximation (since it would give rise to measure zero events).
Appendix H Inference
Algorithm 3 DyDM sampling
1:Input: dimension $n$, schedule $\mathcal{T}=\{t_{j}\}$, trained score network $s_{\theta}$
[…]
4:while $t>0$ do
5:  Let $u\sim\mathcal{N}(0,I_{n})$
6:  $\delta t\leftarrow\text{BackwardStepsizeController}(\lambda(t),u)$ $\triangleright$ Step size as described F.2.
7:  $\hat{s}_{\theta}\leftarrow$ Interpolation of $s_{\theta}(\lambda,t_{j})$ and $s_{\theta}(\lambda,t_{j+1})$ for $t_{j},t_{j+1}\in\mathcal{T}$ closest points in schedule
8:  $\rm BackwardsDrift$$\leftarrow$ drift of eq. 5 at time $t$ and point $\lambda(t)$ with score $s_{\theta}(\lambda,t_{j+1})$
9:  if $\lambda_{t}-{\rm BackwardsDrift}\cdot\delta t-u\sqrt{2\alpha\delta t}\in{\rm Weyl\ Chamber}$ then
10:   $\lambda_{t-\delta t}\leftarrow\lambda_{t}-{\rm BackwardsDrift}\cdot\delta t-u\sqrt{2\alpha\delta t}$
[…]
12:   $\lambda_{t-\delta t}\leftarrow\lambda_{t}-(-{\rm ForwardDrift})\cdot\delta t-u\sqrt{2\alpha\delta t}$ $\triangleright$ $\rm ForwardDrift$} as in 3.1
13:  end if
14:  $t\leftarrow t-\delta t$
15:end while
16:Output: spectral sample $\lambda(0)\in\mathbb{R}^{n}$
Figure 6: Sampling from the Dyson Diffusion Model. The eq. 5 is evolved backward in time with the shooting mechanism and an adaptive step size, ensuring that the paths remain in the Weyl Chamber.
We describe in Algorithm 3 the sampling procedure with the shooting mechanism (see Appendix G for details) incorporated, ensuring that we remain in the Weyl Chamber.
Appendix I Engineering
We follow Karras et al. (2024) by using SiLU activations. We further use EMA to average over multiple runs (Song & Ermon, 2020).
A key strength of DyDM is that the score $s_{\theta}(\lambda,t)$can be parameterized with any learning architecture, without being constrained to GNNs or graph transformers. We demonstrate this by parameterizing the score with a simple MLP. The size of layers varies by application, and we document it for each dataset in Github . For instance, for the large dataset of $15^{\prime}000$ brain ego graphs, our model consists of a hidden MLP of depth $4$, where the input and output layer have width $64$ and the hidden layers have width $256$. The space + time data is first scaled up with a linear layer from size $n+1=11$ followed by a batch norm to feed into the hidden MLP, and upon processing through the hidden MLP, it gets scaled down through a simple linear layer to size $n=10$. In the hidden MLP, we employ as nonlinearities scaled SiLU functions, as argued by Karras et al. (2024).
To make full use of the GPU memory, in one epoch, we sample $N_{paths}$ paths in parallel. If the dataset is too small to fill the GPU memory, we sample multiple, independent, paths of the same data points in parallel. From these samples, we update the score network $s_{\theta}$ in smaller batches. Proceeding in this way ensures that JAX uses the full potential of the GPU. These parameters ($N_{paths}$ and batch size) can be specified in the configuration file.
We implement the model in Jax and Equinox (Kidger & Garcia, 2021).
I.1 Choice of Time Grid
As outlined in Section 3.2, we choose an exponential time grid on which the objective is learned. This is due to the mixing behavior of Dyson’s Brownian Motion. For instance, for the Brain dataset, we use exponential time grid detailed in Table 2.
Table 2: Example of exponential time grid, here for the brain data set which contains in total $15^{\prime}000$ graphs. $\mathrm{d}t$ is $0.05$, and the final time is $T=12.0$.
from to stepsize
0 $1/8$ $1/64\cdot\mathrm{d}t$
$1/8$ $1/4$ $1/32\cdot\mathrm{d}t$
$1/4$ $1/2$ $1/16\cdot\mathrm{d}t$
$1/2$ 11 $1/8\cdot\mathrm{d}t$
11 22 $1/4\cdot\mathrm{d}t$
22 33 $1/2\cdot\mathrm{d}t$
33 77 $1\cdot\mathrm{d}t$
77 T $2\cdot\mathrm{d}t$
I.2 Preprocessing and dealing with different dimensions
As is the case for an Ornstein-Uhlenbeck diffusion, the speed of convergence depends for Dyson-BM on $(1)$ the coefficients and $(2)$ the initial condition. We can thus choose to significantly vary the final time $T$ or rescale the initial condition. We choose the latter (although both options are feasible). To that end, we rescale the spectra with an affine transformation in a preprocessing step. In more detail: For a given set of graphs, we perform an eigendecomposition, and rescale so that among the the entire dataset the largest eigenvalue is at most $\lambda_{\rm max}$ and at least $\lambda_{\rm min}$. For instance, on the benchmark models we chose $\lambda_{\rm max}=5$, $\lambda_{\rm min}=-5$. If a spectrum has eigenvalues of multiplicities greater than $1$, we perform an $\epsilon$-perturbation, where the $\epsilon$ depends on the distance of the closest eigenvalues in the dataset. In postprocessing, the spectra are scaled back and the perturbation is undone for eigenvalues that are at most $\epsilon$ apart after generation. Note that with this preprocessing, only one eigendecomposition per training graph is necessary.
Appendix J Challenges of Dyson’s Brownian Motion
Using Dyson’s Brownian Motion for a diffusion presents several challenges, all of which we overcame in this paper. First, the Dyson SDE is not an OU process, but instead an SDE with singularities of order $\mathcal{O}(1/(\lambda_{k}-\lambda_{l}))$ in the drift, posing both theoretical and numerical challenges. Second, the conditional density $p(x\mid x_{0})$ is non-Gaussian and challenging to obtain, as with any non-OU diffusion process. Therefore, we do not have access to a canonical loss function. Finally, the non-availability of conditional distributions means that training is not simulation-free.
We overcome the obstacles mentioned above and provide a diffusion model for the spectra of graphs based on the Dyson SDE (Fig. 1). The model is not only efficient but is also able to distinguish between spectra of graphs that GNNs are blind to (Fig. 2). In addition, with DyDM, no ad hoc data augmentation is necessary. Further, through Eigenvector-SDE, we give the dynamics of the remaining information in form of conditioned eigenvector dynamics, hence making them accessible for future work devoted to eigenvector diffusion.
J.1 Why not log-transform the SDE?
One idea could be to transform the spectral SDE into terms $\lambda_{1},\lambda_{2}-\lambda_{3},\ldots,\lambda_{n-1}-\lambda_{n}$ and take logarithms, to avoid singularities. However, this is no desirable for multiple reasons: First, upon applying Ito to this SDE, we (i) lose the $\log$ and obtain singularities again and (ii) get higher order singularities $\mathrm{d}\log(x_{t})=\frac{1}{x_{t}}\mathrm{d}x_{t}-\frac{1}{2x_{t}^{2}}(\mathrm{d}x_{t})^{2}$. Second, the space that would need to be sampled would certainly not decrease, since now the transformed domain reaches from -$\infty$ to $\infty$. Hence, we choose the method described in the main part.
Appendix K Why not learn on all $n!$ many graph representations?
In short, learning $n!$ more data is much harder. This point has been mentioned by previous literature, e.g. Niu et al. (2020). However, if we go deeper, the interested reader might wonder why exactly.
K.1 Rigorous argument
We give here a toy example, where the challenge can be phrased rigorously. Suppose we have a binary matrix $X\in\{0,1\}^{m,k}$ with independent entries, which are for $i\in[m],j\in[k]$ distributed as $X_{ij}\sim\operatorname{Bernoulli}\left(p_{j}\right)$ for some unknown $p_{j}\in[0,1]$. The task is to estimate the $p_{j}$. The motivation for this example stems from the following setting: We want to learn the probability of $k$ objects (for instance, graphs), each having $m$ representations (for instance, representations of graphs such as adjacency matrices). Each column of the matrix $X$ thus consists of all representations of the same object. We define two estimators, with $\rm est_{A}$} using the inductive bias and $\rm est_{B}$} not using it.
To that end, suppose we have $N$ uniformly at random obtained samples $Z_{1},\ldots,Z_{N}$. In more detail, that means that we sample for each $\ell\in[N]$ a pair of indices $(i_{\ell},j_{\ell})\in[m]\times[k]$ uniformly at random, and describe the obtained sample by $Z_{\ell}\sim X_{i_{\ell},j_{\ell}}$. Estimator $\rm est_{A}$} makes use of the inductive bias. That is, for $u\in[m],v\in[k]$ we define
${\rm est_{A}^{(u,v)}}\coloneq\frac{k}{N}\sum_{\ell\in[N]}Z_{\ell}\mathds{1}_{j_{\ell}=v}.$
Estimator $\rm est_{B}$} does not make use of the inductive bias. That is, for $u\in[m],v\in[k]$, we define
${\rm est_{B}^{(u,v)}}\coloneq\frac{k\cdot m}{N}\sum_{\ell\in[N]}Z_{\ell}\mathds{1}_{i_{\ell}=u,j_{\ell}=v}.$
Clearly, both estimators are unbiased: $\operatorname{\mathbb{E}}\left[{\rm est_{A}^{(u,v)}}\right]=\operatorname{\mathbb{E}}\left[{\rm est_{B}^{(u,v)}}\right]=p_{v}$. However, their mean squared error, defined for $X\in\{A,B\}$ as
${\rm MSE}({\rm est_{X}})\coloneqq\frac{1}{m\cdot k}\sum_{u\in[m],v\in[k]}\operatorname{\mathbb{E}}\left[\left({\rm est_{X}^{(u,v)}}-p_{v}\right)^{2}\right],$
varies significantly between $\rm est_{A}$} and $\rm est_{B}$}.
Corollary K.1 (Mean square error).
In dependence of the problem size $m$ and number of samples $N$, the mean square error of $\rm est_{A}$} is of order $\Theta(1/N)$, while the mean square error of $\rm est_{B}$} is of order $\Theta(m/N)$.
To prove Corollary K.1, we derive the mean squared error for both estimators.
Lemma K.2 (MSE of $\rm est_{A}$}).
The mean squared error of estimator $\rm est_{A}$} is
${\rm MSE}\left({\rm est_{A}}\right)=\frac{1}{N}\sum_{v\in[k]}\left(p_{v}-\frac{1}{k}p_{v}^{2}\right).$
Proof.
We have for $u\in[m],v\in[k]$
$\operatorname{\mathbb{E}}\left[\left({\rm est_{A}^{(u,v)}}-p_{v}\right)^{2}\right]$
$=$ $\operatorname{\mathbb{E}}\left[\left(\frac{k}{N}\sum_{\ell\in[N]}Z_{\ell}\mathds{1}_{j_{\ell}=v}-p_{v}\right)^{2}\right]$
$=$ $\frac{k^{2}}{N^{2}}\sum_{\ell\in[N]}\operatorname{\mathbb{E}}\left[\sum_{o\in[N]}Z_{\ell}\mathds{1}_{j_{\ell}=v}X_{o}\mathds{1}_{j_{o}=v}\right]-2\frac{k}{N}\operatorname{\mathbb{E}}\left[\sum_{\ell\in[N]}Z_{\ell}\mathds{1}_{j_{\ell}=v}\right]p_{v}+p_{v}^{2}$
$=$ $\left(1-\frac{1}{N}\right)p_{v}^{2}+\frac{k}{N}p_{v}-2p_{v}^{2}+p_{v}^{2}$
$=$ $\frac{k}{N}p_{v}-\frac{1}{N}p_{v}^{2}.$
Averaging over all $u\in[m],v\in[k]$, we obtain the desired result. ∎
Lemma K.3 (MSE of $\rm est_{B}$}).
The mean squared error of estimators $\rm est_{B}^{(u,v)}$} for $u\in[m]$, $v\in[k]$ is
${\rm MSE}\left({\rm est_{B}}\right)=\frac{1}{N}\sum_{v\in[k]}\left(m\cdot p_{v}-\frac{1}{k}p_{v}^{2}\right).$
Proof.
We have for $u\in[m],v\in[k]$
$\operatorname{\mathbb{E}}\left[\left({\rm est_{B}^{(u,v)}}-p_{v}\right)^{2}\right]$
$=$ $\operatorname{\mathbb{E}}\left[\left(\frac{k\cdot m}{N}\sum_{\ell\in[N]}Z_{\ell}\mathds{1}_{i_{\ell}=u,j_{\ell}=v}-p_{v}\right)^{2}\right]$
$=$ $\frac{k^{2}m^{2}}{N^{2}}\sum_{\ell\in[N]}\operatorname{\mathbb{E}}\left[\sum_{o\in[N]}Z_{\ell}\mathds{1}_{i_{\ell}=u,j_{\ell}=v}Z_{o}\mathds{1}_{i_{o}=u,j_{o}=v}\right]-2\frac{k\>m}{N}\operatorname{\mathbb{E}}\left[\sum_{\ell\in[N]}Z_{\ell}\mathds{1}_{i_{\ell}=u,j_{\ell}=v}\right]p_{v}+p_{v}^{2}$
$=$ $\left(1-\frac{1}{N}\right)p_{v}^{2}+\frac{k\>m}{N}p_{v}-2p_{v}^{2}+p_{v}^{2}$
$=$ $\frac{k\>m}{N}p_{v}-\frac{1}{N}p_{v}^{2}$
Summing over all estimators for $u\in[m],v\in[k]$ gives the desired result. ∎
By considering the estimation problem as a problem of parameters $m$, NN, Corollary K.1 follows directly from Lemma K.2, Lemma K.3.
The same argument may be carried out with Normal instead of Bernoulli random variables.
Appendix L WL-equivalence of regular graphs
We now prove Lemma 2.1.
Lemma 2.1 (restated).
For every fixed $n,k\in\mathbb{N}$, all $k$-regular graphs $G\in\mathcal{G}^{n}$ are WL equivalent. Moreover, every graph $G\in\mathcal{G}^{n}$ that is WL equivalent to a $k$-regular graph is $k$-regular.
Proof.
Recall that the 1-Weisfeiler-Leman (WL) algorithm tests graph equivalence by iteratively updating vertex colors. Initially, all vertices share one color. In each step, a vertex’s new color is determined by the multiset (a set allowing for duplicates) of its own color and its neighbors’ current colors. This continues until the coloring stabilizes. Two graphs are WL-equivalent if this process generates identical color counts (histograms) at every step.
For a $k$-regular graph, since every vertex has exactly $k$-neighbors, in each iteration of the WL-algorithm all vertices have the same color. Thus in particular the histograms are always the same and therefore any two $k$-regular graphs are WL-equivalent.
A $k$-regular graph is not WL-equivalent to a $\ell$-regular graph for $k\neq\ell$ since the colors after the first iteration are distinct as the number of neighbors is different. Moreover, if a graph is not regular, then the first iteration must assign a different new color to at least two vertices in the first iteration. Therefore, the color histogram is not the same as the color histogram of a regular graph and thus they are not WL-equivalent. This shows that the WL-equivalence class of a $k$-regular graph is exactly the set $k$-regular graphs. ∎
In particular, it follows from Morris (Morris et al., 2019) that GNNs cannot distinguish $k$-regular graphs.
Appendix M Datasets
WL-bimodal
The WL-Bimodal graph consists of $80\%$ graph $A$ and $20\%$ graph B (see Fig. 2) adjacency matrices. We drew among all permutations $5^{\prime}000$ permutations uniformly at random and shuffled the graphs. The first $80\%$ of this dataset are used for training, the remaining $20\%$ are used for testing.
Community-small
This standard benchmark dataset (Niu et al., 2020; Jo et al., 2022; You et al., 2018) consists of $100$ graphs of size up to $20$ vertices. We comment in Section N.1 on the small dataset set ($100$ graphs) compared to the big dimension (up to 20 vertices) and the thereby induced effect of undersampling.
Brain
We report this dataset in our repository. In detail, we construct from the brain graph Amunts et al. (2013); Rossi & Ahmed (2015) so-called ego-graphs. That is, we take the (distance 1) neighborhoods of vertices, and consider the induced subgraph. From those, we generate $15^{\prime}000$ graphs of size $n=5$ to $n=10$ vertices with eigenvector multiplicity up to $3$, with the closest eigenvalues – which are not multiplicities – having distance $0.036$. We take $70\%$ as train graphs, $15\%$ as validation, and the remaining $15\%$ as test graphs.
Appendix N Comparing to Benchmark Models
N.1 On undersampling
For the “bimodal” case, we have sufficient statistics for the $10$-dimensional space $C_{10}$ ($N=5^{\prime}000$ graph samples, each isomorphic to one of two graphs) and know in addition the underlying distribution; hence, an extensive interpretation of this result is appropriate. For a fair comparison, we thus follow the standard test/train split procedure as reported in (Jo et al., 2022; You et al., 2018; Niu et al., 2020) using $80\%$ of the data as train data and the remaining $20\%$ as test data.
Conversely, the standard benchmark set “community small” (Niu et al., 2020; Jo et al., 2022; You et al., 2018) contains only 100 graphs, and each has a size of up to $n=20$ vertices. Thus, a comparison from the learned distribution based on (few) training samples to (very few) test samples suffers from undersampling. This becomes very stark if one considers the following issue: If one would extract $80\%$ of the dataset for training, where they are taken from already matters a lot: Whether they are taken from the front or back changes the maximal graph size in the training set. Depending on whether the training data is taken from the front or back of the dataset, a model trained on the training set might thus have no possibility to learn the correct maximal graph size. More generally, poorness of benchmarks in graph generative learning has been recently addressed by Bechler-Speicher et al. (2025). To offer some consistent comparison, we do include the standard benchmark “community small”, but focus on memorization rather than the (on those benchmarks untestable) generalization.
To overcome the issue of undersampling, we construct a set of $15^{\prime}000$ ego-graphs from the brain dataset Amunts et al. (2013) as described in Appendix M, which is sufficiently large to not suffer from undersampling. We train both our model and DiGress on $70\%$ (= $10^{\prime}500$ graphs), perform hyperparameter tuning (see below for details on the DiGress hyperparameter tuning) on a validation set of $15\%$, and test on the remaining $15\%$.
N.2 GDSS
To compare to the GDSS model Jo et al. (2022), we take the following approach to ensure a fair comparison: The GDSS model has been trained and optimised on the ego small and community small dataset, so that we take for these datasets the snapshots and hyperparameters given by the original paper Jo et al. (2022). For the remaining datasets, we start with the settings from the community-small dataset since that has similar size, adapt the maximal number of vertices to the dataset, and then perform hyperparameter tuning as described in Appendix C of the GDSS paper: We form a grid search on the model’s following parameters: The scale coefficient in $\{0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,1.0\}$, the signal-to-noise ratio in $\{0.05,0.1,0.15,0.2\}$, and in addition to the GDSS paper, we also try different $\beta_{max}$ in $\{1,10,20\}$, and try different batch sizes $\{128,4096\}$. We test with and without EMA. The motivation for the additional parameters we hyper tune is that we observed that they further improve the GDSS model. This gives full fairness to the model. The training time on one H100 GPU per job was about 24 hours for each of the configurations with batch size $128$ and approximately 1 hour and 45 minutes for each of the configurations with batch size $4096$.
On the Two-WL graph case, our hyperparameter tuning resulted in $280$ models that we trained and sampled from. From each of those 280 models, we generated $1^{\prime}000$ graphs. Most of those models worked fine, that is 269 models did not contain NaNs in their output. From those, we select the best model based on the following relative error: From the generated samples, we calculate the share of spectra $\epsilon$-close to the spectrum of graph $A$, say $\hat{p}_{A}$ and the share of spectra $\epsilon$-close to graph $\hat{p}_{B}$ (we choose the $l_{2}$ distance with $\epsilon=0.2$). Recall that in the ground truth, we have $p_{A}=0.8$, $p_{B}=0.2$. We then selected the best model based on the relative error
[…]
In summary, we invested a lot of resources in following both the hyperparameter tuning given in the GDSS paper and, in addition, tried new hyperparameters, leading to the $280$ models that we trained and sampled from. This ensures maximal fairness.
N.3 EDP-GNN
We proceeded analogously to GDSS: We used the given configurations for community small. We observe that the model already performs hyperparameter tuning of the noise scales during sampling. For the other data sets, we have used the given configurations for community small and in addition tried the learning rates $\{0.001,0.0002\}$, number of diffusion steps $\{1^{\prime}000,2^{\prime}000\}$ and number of layers $\{4,6\}$. In the two graph case, for example, the optimal configuration was with learning rate $0.0002$, number of diffusion steps being equal to $2^{\prime}000$ and $6$ layers. The training time for $5^{\prime}000$ epochs on a H100 GPU was approximately 10 hours.
N.4 DiGress
As with the previous models, we used the given configurations for community small. For the other data sets, we have used the given configurations for community small and in addition tried the learning rates $\{0.001,0.0002\}$, weight decay parameters $\{10^{-2},10^{-12}\}$, number of diffusion steps $\{500,1^{\prime}000\}$ and number of layers $\{5,8\}$. The model was quite robust to hyperparameter tuning and in the two graph example, for all parameters the model sampled the $80\%$-graph with likelihood between $75\%$ and $82\%$. The training time for $1^{\prime}000$ epochs on a H100 GPU was approximately 3 hours.
N.5 Congress
Since Congress and DiGress come from the same paper, we have proceeded almost exactly as in Digress. We tried the learning rates $\{0.001,0.0002\}$, weight decay parameters $\{10^{-2},10^{-12}\}$, number of diffusion steps $\{500,1^{\prime}000\}$ and number of layers $\{6,8\}$. The model was again rather robust to hyperparameter tuning, yet not as much as Digress. In the two graph example, for all parameters the model sampled the $80\%$-graph with likelihood between $10\%$ and $25\%$. The default learning rate from the community-small configuration was $0.0002$, yet we have observed that the results were significantly better with learning rate $0.001$. The training time for $1^{\prime}000$ epochs on a H100 GPU was approximately 3 hours.
N.6 Comparison table with more digits
We provide here the table shown in the main part Table 1 but with more digits (not implying that all are statistically significant): This rationalizes which entries in Table 1 are dark green and which ones are light green. Note that this result is included only for transparency, since the reported digits here go beyond the significant digits.
Table 3: Statistical distances of DyDM compared to standard models, as in Table 1 but with more digits.
Dataset
WL-Bimodal
Community Small Brain
Distance $\mu$ $\mathcal{W}_{\mathrm{marg}}$ $\mu$ $\mathcal{W}_{\mathrm{marg}}$ $\mu$ $\mathcal{W}_{\mathrm{marg}}$
DyDM (ours) $\bm{0.0166}$ $\bm{0.0076}$ $\bm{0.0671}$ $\bm{0.0172}$ $\bm{0.0455}$ $\bm{0.0275}$
EDP-GNN 0.1342 0.0750 0.4164 0.1356 - -
GDSS 0.2289 0.1252 0.4180 0.1444 - -
ConGress 0.3802 0.1590 0.2741 0.1138 - -
DiGress (no trick) 1.0568 0.2852 2.5088 0.4481 - -
DiGress (trick) 0.0302 $\bm{0.0073}$ 0.0934 0.0254 0.1208 $\bm{0.0285}$
Appendix O Learning dynamics of EDP-GNN
We report in Figure 7 the learning progress of EDP-GNN on the WL-bimodal dataset. We average over $4$ different training and sampling runs of EDP-GNN. After $5^{\prime}000$ epochs, we observe the result of EDP-GNN reported in Figure 2.
We observe that the model quickly learns the WL equivalence class (the pink line is from epoch $500$ onward close to $1$). The share of graph $A$ and graph $B$ samples initially increase until epoch $1^{\prime}500$, but then remain low and significantly different from the ground truth (blue and green dashed lines). Importantly, a significant share of the samples are $WL$-equivalent but neither isomorphic to graph $A$ nor to graph $B$.
Figure 7: Learning dynamics of EDP-GNN on the WL-bimodal dataset (ground truth = $80\%$ graph A, $20\%$ graph B): The model learns very quickly (after less than 500 epochs) the WL-equivalence class, but struggles to learn graphs $A$ and $B$.